### v1

# Holistic Optimization Framework for FPGA Accelerators

Abstract
Customized accelerators have transformed modern computing by enhancing energy efficiency and performance through specialization. Field Programmable Gate Arrays play a pivotal role in this domain due to their flexibility and high-performance potential. High-Level Synthesis and source-to-source compilers simplify hardware design by converting high-level code into hardware descriptions enriched with directives. However, achieving high Quality of Results in FPGA designs remains challenging, requiring complex transformations, strategic directive use, and efficient data management. While existing approaches like Design Space Exploration (DSE) and source-to-source compilers have made strides in improving performance, they often address isolated aspects of the design process.
This paper introduces Prometheus, a holistic framework that integrates task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution into a unified design space. Leveraging Non-Linear Problem methodologies, Prometheus explores this space to find solutions under resource constraints, enabling bitstream generation.
I Introduction
The emergence of customized accelerators has revolutionized modern computing by delivering significant energy efficiency and performance gains through specialization. Field-Programmable Gate Arrays (FPGAs) have gained prominence for their flexibility and high-performance potential. High-level synthesis (HLS) and source-to-source compilers [19, 81, 76, 32, 62, 50, 77, 31, 79, 38, 80, 43] have made hardware design more accessible by converting high-level code (e.g., C++, Python) with hardware-specific directives into hardware descriptions. However, their effectiveness heavily relies on the quality of the input code, the applied hardware directives, and the use of advanced code transformations and directive strategies to optimize data communication and achieve high Quality of Result (QoR).
Challenges in Existing Approaches  Several works have contributed to enhancing performance through Design Space Exploration (DSE) [58, 59, 67, 66, 64, 24, 82, 66] and source-to-source compilers [38, 79, 78, 60, 71]. Despite these advancements, significant challenges remain, notably in supporting Code Transformation, which involves efficiently optimizing code to fully exploit FPGA potential; Overlapping Communication and Computation, aimed at maximizing concurrency to enhance performance; Concurrent Task Execution, which ensures simultaneous execution of tasks without resource conflicts; and Exploiting Intra-Task Parallelism, leveraging parallel execution within tasks to further boost performance. Previous works address specific optimization challenges but do not tackle all aspects simultaneously, underscoring the need for comprehensive, integrated solutions. Moreover, generating synthesizable code that can be converted into a bitstream without human intervention is a persistent challenge in FPGA design. This complexity is exacerbated by the intricate architecture of FPGAs, which often incorporates multiple Super Logic Regions (SLRs). In addition to adhering to the overall resource constraints of the entire FPGA board, it is crucial to meet specific constraints of each individual SLR while the following factors need to be considered to generate a high-quality HLS solution.
Shared Buffer  A classical approach to improving performance involves using shared buffers, which cache data on-chip to facilitate reuse by various computational tasks. This method increases parallelism and is particularly effective for computation-bound kernels. Studies have shown the effectiveness of shared buffers [66, 64, 79, 77, 38, 31, 60, 59, 58]. However, on-chip memory limitations often prevent caching all necessary data, and overlapping computation and communication remains constrained. While some compilers like Merlin [77] and PASTA [35] support double buffering for single arrays, Merlin’s application is limited and context-specific, and PASTA lacks parallelism.
Dataflow  On FPGAs, dataflow enables seamless data movement between off-chip and on-chip memory and computational tasks. This method supports the overlap of communication and computation, enabling concurrent execution of tasks and allowing dependent tasks to proceed once data is available via FIFO (First-In-First-Out). Dataflow is highly effective for memory-bound kernels but limits parallelism due to the absence of shared memory among tasks. Each task is compiled independently, which simplifies the compilation process and can enhance operating frequency [26, 27, 28].
Integrating shared buffers with dataflow methods combines their strengths, boosting performance and reducing resource use. A specialized design space for automated bitstream synthesis streamlines the process by reducing manual rewriting of non-synthesizable code, requiring efficient global and per-task resource management.
Prometheus Approach  To address these limitations and advance high-level source-to-source compilation, we propose Prometheus. This framework integrates the benefits of shared buffers and dataflow, aiming to improve performance while managing resources efficiently to ensure bitstream generation. The main contributions are:
  • The introduction of a unified design space facilitates the development of a tiled dataflow architecture. This space incorporates task fusion, tiling strategies, loop permutation, overlapping of computation and communication automatically, and concurrent task execution.
  • Employing Non-Linear Problem methodology to navigate and discover theoretically optimal solutions within stringent resource constraints for generating bitstreams.
  • Automating our framework that is end-to-end comprehensive. It automatically generates a host file in OpenCL to verify correctness, runs Vitis HLS for report generation and RTL simulations, and generates the bitstream.
  • Conducting evaluations utilizing RTL simulation and on-board FPGA runs, delivering superior Quality of Results (QoR) compared to AutoDSE [67], Sisyphus [60], ScaleHLS [79], and Allo [14]. Importantly, it achieves these results without requiring DSE for HLS design generation.
II Background and Motivation
II-A HLS optimization
In High-Level Synthesis (HLS), directives such as unroll, pipeline, and array_partitioning pragmas play a crucial role. The unroll pragma enables parallel execution by duplicating statements, while pipeline stages loops for overlapping operations based on the initiation interval (II). array_partitioning divides arrays into on-chip memory banks, facilitating simultaneous data access and enhancing unrolling. Additional optimizations include shared buffers for reducing off-chip transfers, coarse-grained parallelism through duplicated processing elements, and ping-pong buffering to overlap data transfer and computation [18]. On AMD/Xilinx FPGAs, data transfers use bit widths up to 512 bits, with the maximum bit width determined by problem dimensions. For instance, transferring 216 floats with a 256-bit width (8 floats per transfer) requires 27 transfers. The dataflow pragma organizes computation into actors that communicate via FIFO queues, where each actor processes inputs and produces outputs as they become available [41].
Performance evaluation methods include estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation. Estimation and simulation avoid place-and-route but utilize board resources, while on-board evaluation imposes stricter constraints for bitstream generation.
II-B Limitation of existing methods
To demonstrate the limitations of the existing framework and methodology we use 3mm. As illustrated in Listing 1, two independent matrix multiplications compute matrices that are then utilized in the final matrix multiplication.
Shared Buffer  AutoDSE [67], Sisyphus [60], ScaleHLS [79], and frameworks like [66, 64, 59, 58] optimize the 3mm kernel by sequentially executing each matrix multiplication. AutoDSE uses a bottleneck-based DSE approach, Sisyphus employs loop splitting and NLP-based cost modeling, and ScaleHLS automates multi-level optimizations in Vitis HLS.
Sequential execution enables DSP reuse but lacks support for concurrent task execution, e.g., the first two matrices in 3mm are independent and can run concurrently. Although arrays can be transferred at different loop levels to manage on-chip memory, automatic computation-communication overlap is not achieved. Current frameworks enhance performance by transferring large arrays on-chip and extensively unrolling loops to maximize parallelism.
Concurrent execution of the first two matrix multiplications could be achieved using a shared buffering method, but this remains unexplored. Extending concurrency to all three operations would require fusion techniques, potentially limiting the compiler’s optimization capabilities.
1for (i = 0; i < 180; i++) // MM 1
2 for (j = 0; j < 190; j++) {
3 E[i][j] = 0.0; // S0
4 for (k = 0; k < 200; ++k)
5 E[i][j] += A[i][k] * B[k][j];} // S1
6for (i = 0; i < 190; i++) // MM 2
7 for (j = 0; j < 210; j++) {
8 F[i][j] = 0.0; // S2
9 for (k = 0; k < 220; ++k)
10 F[i][j] += C[i][k] * D[k][j];} // S3
11for (i = 0; i < 180; i++) // MM 3
12 for (j = 0; j < 210; j++) {
13 G[i][j] = 0.0; // S4
14 for (k = 0; k < 190; ++k)
15 G[i][j] += E[i][k] * F[k][j];} // S5
Listing 1: Code of 3mm
Dataflow  The dataflow method creates six tasks, each corresponding to a statement, with communication managed through FIFOs. For the 3mm kernel, efficient data reuse is critical and can be achieved via reuse buffers or repeated off-chip data transfers. However, reuse buffers are limited to individual tasks and cannot be shared across tasks.
As 3mm is computation-bound, increasing parallelism is vital. However, parallelism is constrained by the receiving task’s FIFO bit width, and attempts to scale this approach often lead to routing congestion and impractical bitstream generation. Despite these challenges, dataflow allows concurrent execution of the first two matrix multiplications and starts the third as soon as data becomes available.
Mixed Method  Methods like Allo [14] and HIDA [78] integrate shared buffers and dataflow techniques to enhance QoR. Allo extends HeteroCL [38] within the MLIR framework [40], supporting code transformation and dataflow but lacks automatic data-tile selection and pragma insertion.
Limitations in General  Existing HLS frameworks often lack key optimizations, such as automatic ping-pong buffering for overlapping computation and communication, tile size adjustments, and padding to maximize bit width for reducing communication latency. These gaps can lead to suboptimal performance, particularly in data-intensive scenarios. Most frameworks [60, 59] restrict unroll factors to evenly divide problem sizes, while others [67, 66] leveraging Merlin [77] permit uneven unroll factors. However, optimizations typically target complete tiles, resulting in inefficiencies with partial tiles and non-innermost unrolled loops. Expanding unroll factor options could better balance parallelism and performance.
For the 3mm kernel optimized with RTL simulation, our framework (Prometheus) achieved a throughput of 368.36 GF/s, significantly surpassing Sisyphus (178.97), Allo (60.40), ScaleHLS (43.04), and AutoDSE (1.74). These frameworks employ varying strategies—AutoDSE, Sisyphus, and ScaleHLS use shared buffer methods, while Allo applies a mixed method. However, due to the limitations noted, further improvements are needed to achieve a superior QoR.
II-C Overview of Prometheus
We propose Prometheus, a unified optimization framework leveraging a Non-Linear Problem (NLP) cost model to explore design spaces efficiently. Prometheus is engineer-friendly, offering adjustable parameters such as tile size, array partitioning, and unroll factors, while being SLR-aware for task distribution across SLRs, simplifying bitstream generation and optimizing performance-resource balance.
The pseudo-code in Listing 2 illustrates how Prometheus processes the 3mm kernel. The load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFOs. Each task_i represents the fully unrolled computation of the intra-tile for statement Si in Listing 1. We can see an example on Listing 3. Arrays E, F, and G are initialized to 0 with statements S0, S2, and S4, respectively, and are not loaded. Prometheus optimizes computation by fusing statements producing the same outputs (e.g., Listing 1), automatically generating an NLP problem to determine loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. It strategically adjusts bit widths (e.g., for arrays F, D, and G) to balance parallelism and resource usage.
1/*************** Fused Task 0 ****************/
2float B[204][192]; load_B(B);
3for (i0 = 0; i0 < 18; i0++)//inter-tile loop
4 float A[10][204]; load_A(A);
5 for (j0 = 0; j0 < 6; j0++){//inter-tile loop
6 float E[10][32];
7 task_0(E); // S0 + intra-tile loops
8 for (k0 = 0; k0 < 51; ++k0)//inter-tile loop
9#pragma HLS pipeline II=3
10 task_1(E, A, B); // S1 + intra-tile loops
11 store_E(E); sent_E(E);
12/*************** Fused Task 1 ****************/
13float C[190][222]; load_C(C);
14for (j0 = 0; j0 < 7; j0++)//inter-tile loop
15 float D[222][32]; load_D(D);
16 for (i0 = 0; i0 < 10; i0++){//inter-tile loop
17 float F[19][32];
18 task_2(F); // S2 + intra-tile loops
19 for (k0 = 0; k0 < 74; ++k0)//inter-tile loop
20#pragma HLS pipeline II=3
21 task_3(F, C, D); // S3 + intra-tile loops
22 store_F(F); sent_F(F);
23/*************** Fused Task 2 ****************/
24float E[180][192];
25for (j0 = 0; j0 < 7; j0++)//inter-tile loop
26 float F[192][32];
27 receive_F(F);
28 for (i0 = 0; i0 < 18; i0++){//inter-tile loop
29 float G[10][32];
30 receive_E(E);
31 task_4(G); // S4 + intra-tile loops
32 for (k0 = 0; k0 < 32; ++k0)//inter-tile loop
33#pragma HLS pipeline II=3
34 task_5(G, E, F); // S5 + intra-tile loops
35 store_G(G);
Listing 2: Pseudo Code of 3mm generated by Prometheus
Efficient overlap of computation and communication is achieved via ping-pong buffering, concurrent task execution, and FIFO-triggered dependent tasks. Fused tasks 0 and 1 execute concurrently, and Task 2 starts once arrays F and E are ready. These optimizations enable Prometheus to surpass frameworks like Sisyphus [60] in throughput and QoR, particularly for computation-bound kernels. Its SLR-awareness and dataflow integration make it practical for generating deployable, resource-efficient bitstreams.
III Problem Formulation
In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time. This compile-time awareness enables us to construct a precise model that facilitates rigorous optimizations. To leverage FPGA parallelism, we implement an acyclic dataflow graph, ensuring parent nodes do not receive data from their children. While this constraint limits graph configurations and may increase resource usage, it reduces overall latency. To support this structure, we inline [15] each function in the input code to generate the required acyclic graph. Our primary objective is to minimize latency within resource constraints by overlapping communication and computation within tasks and executing independent tasks concurrently. To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section IV.
Dependency Graph Creation  Our process begins with the input C/C++ code, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC [69] verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, where each loop body corresponds to a single task, we construct a dependency graph capturing inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
Prometheus’s solution space includes:
Data-tiling and Padding  From the fused dependency graph, edges represent data communication between tasks and off-chip memory, transferring data-tiles of specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. If an array is not fully transferred in a dimension $n$, its tile size is constrained by other arrays sharing the same loop. To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. For example, in Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting the outermost loop. Using ISCC [69], we ensure the legality of permutations, yielding two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is infeasible, the inter-tile loop remains in its legal position. In Listing 2, inter-tile loops (lines 3 and 5) iterate intra-tile loops in task_0, which are fully unrolled to maximize parallelism. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$. Fused tasks are output-stationary, ensuring each tile’s output is processed (loaded, computed, and stored or sent) only once. Dependent tasks execute sequentially with high throughput, leveraging unrolled tasks and on-chip data. Processed data-tiles are then transferred to the next task or off-chip memory, optimizing performance.
Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP, as described in Section IV. As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
Automatic Overlapping of Communication and Computation  To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section IV). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only arrays, and triple-buffering for arrays that are both read and written. Following [18], we perform an initial load, then overlap loading the next tile with computing the current one.
Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks  Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution. This concurrency significantly increases overall performance.
Memory Transfer  Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory communication.
IV NLP Formulation
To determine the tile size, loop order, bit width, and memory transfers, we formulate a cost model as a NLP problem aimed at minimizing overall latency. This approach builds upon the methodology proposed by Pouget et al. [59, 58, 60], which we have adapted to address our specific requirements and constraints. In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section III. We employ PoCC [54] to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement. ISCC [69] then generates all legal permutations for each loop body.
Constant Description
$II_{l}$ II of the loop $l$
$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
$TC^{l}_{ori}$ Original Trip Count of the loop $l$
$f_{a,l}$ Footprint of the array $a$ if transferred to on-chip after the loop $l$
$N_{FT}$ Number of fused task
$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation $op$
$DSP$ Number of DSP available for the FPGA used
$max_{part}$ Maximum array partitioning
$SLR$ Number of SLR available for the FPGA used
Variable Description
$TC^{l}_{intra}$, $TC^{l}_{inter}$ TC of the loop $l$ for the intra and inter tile
$TC^{l}$ Trip Count of the loop $l$ after padding
$S^{last}_{a}$ Size of the last dimension of the array $a$ transferred on-chip
$BW_{a}$ Bit width of the array $a$
$t_{a,l},d_{a,l}$ Boolean to know if the array $a$ is transferred and defined (respectively) on-chip after the loop $l$ in the inter-tile
$p_{i}^{l}$ Position of the permutation $i$ for the loop $l$
$slr_{t}$ ID of the SLR use by the task $t$
Set Description
$\mathcal{L},\mathcal{A},\mathcal{S}$ The set of loops, arrays and statements
$\mathcal{L}_{s}$ The set of loops which iterate the statement $s$
$\mathcal{L}_{s}^{red}$ The set of reduction loops which iterate the statement $s$
$\mathcal{L}_{a}^{last}$ The set of loop which iterate the last dimension of the array $a$
$\mathcal{L}_{inter},$ $\mathcal{L}_{intra}$ The set of loops which belong to the inter-tile and intra-tile respectively
$\mathcal{B}$ Set of possible burst size for the data type
$\mathcal{C}_{a_{d}}$ The set of loops which iterates the array $a$ at the dimension $d$
$AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
$\mathcal{P}_{s}$ All permutation of the loops which iterate the statement $s$
$\mathcal{F}_{i}$ The set of statements which belong to the fused task $i$
$\mathcal{T}$ The set of tasks in the dataflow
TABLE I: Overview of the constant, variable and set employed in formulating the NLP
IV-A Constraints
We now describe the constraints by using the code of Listings 1, 2 and 3.
1for (int j1 = 0; j1 < 32; j1++) {
2#pragma HLS unroll
3 for (int i1 = 0; i1 < 19; i1++) {
4#pragma HLS unroll
5 for (int k1 = 0; k1 < 3; k1++) {
6#pragma HLS unroll
7 j=j0*32+j1; i=i0*19+i1; k=k0*3+k1;
8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 3: Pseudo Code of the task_3
Data-tiling and Unroll Factor  The intra-tile transformation, as explained in Section III, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 3, for the array $F$, the loop $j$ (line 7 in Listing 1) has been split into $j0$ (line 14 in Listing 2) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
$\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$ (1)
$(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$ (2)
Bit Width  $\mathcal{B}$ represents the possible bit width for the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
$\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$ (3)
Permutation  Equation 4 ensures that the NLP selects a unique permutation for each loop that iterates through every statement. Equation 5 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
$\forall s\in\mathcal{S},\forall i\in\mathcal{P}_{s},\sum_{p}p_{i}^{l}=1$ (4)
$\forall i\in\llbracket 0,N_{FT}\llbracket,\forall(ft_{0},ft_{1})\in\mathcal{F}_{i}^{2},$
$\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$ (5)
Transfer and Reuse  Equation 6 permits the selection of a single level where each array can be defined (and reused) and transferred. Equation 7 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
The array $E$, defined on line 24 in Listing 2, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 7 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 2, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
$\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$ (6)
$\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$ (7)
On-chip Memory  Equation 8 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
$\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$ (8)
with $N_{a}$ being the number of double buffer for the array $a$.
Array Partitioning  Equation 9 limits the maximum partitioning of each array. This partitioning is crucial as it impacts the maximum unroll factor, necessitating the distribution of data across different BRAM banks under fully unrolled loops, thereby influencing the utilization of BRAMs. Equation 10 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
Array D in Listing 3 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
$\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$ (9)
$\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$ (10)
DSP Utilization  Equation 11 constrains the number of DSPs used based on the available DSP resources. In contrast to [59, 58, 60], we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously.
Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
$\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$ (11)
SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 12). Additionally, Equations 8 and 11 are applied to each SLR to manage resource allocation per SLR effectively.
$\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$ (12)
IV-B Objective Function
The objective function models the dependency graph of the fused task. Independent tasks execute simultaneously, with latency set by the longest task. Dependent tasks overlap partially, with the second task starting once it receives enough data from the first. The latency is the maximum of both tasks’ durations plus the “shift”, the interval before the second task starts. For our 3mm example, the graph is shown in Figure 1, and the corresponding latency is:
$\left.\begin{cases}&Lat(T_{0},T_{1})=\max(Lat(T_{0}),Lat(T_{1})+shift_{T_{0},T_{1}})\\ &Lat(T_{2},T_{3})=\max(Lat(T_{2}),Lat(T_{3})+shift_{T_{2},T_{3}})\\ &Lat(T_{0},..,T_{4})=\max(Lat(T_{0},T_{1}),Lat(T_{2},T_{3}),Lat(T_{4}))\\ &Lat=\max(Lat(T_{0},..,T_{4}),Lat(T_{5})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\max(shift_{T_{1},T_{5}},shift_{T_{3},T_{5}},shift_{T_{4},T_{5}}))\\ \end{cases}\right.$
T0T1T2T3T4T5
Figure 1: Dataflow graph of 3mm maximally distributed
The latency of the intra-task for a statement $s$ is determine similarly to [59, 58, 60] with Equation 13. And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 14.
$Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$ (13)
$Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$ (14)
For each level $n\geq 0$ (with the outermost level being 0) of the inter-tile, we consider the shifting of load, compute, and store operations to enable overlapping of computation and communication. The overlap factor $\alpha$ is set to 1 for read-only operations and 2 for read and store operations:
$Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$ (15)
And then, the final objective function is computed using the dataflow graph, where the latency of the task is established by the latency of the outermost level, denoted as $Lat_{0}$.
V Code Generation
Prometheus takes as input affine C/C++ code and automatically produces an HLS-C++ file, OpenCL host files, and all necessary files for code verification, RTL simulation, and bitstream generation. The NLP described in Section IV gives the parameters of the space such as loop order, tiling factor, etc. However, in order to be efficient, the code generation needs some specific rules.
Communication with Off-Chip Memory  To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler. Furthermore, to guarantee that the load operation transfers data with the correct bit width, we automatically restructure the data in off-chip memory to enable sequential loading. Once data accumulates in the FIFO, a read function moves it to the shared data-tile buffer, whose size is determined by the NLP. For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
Communication between the Fused Task  The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
Intra-Task  Each intra-task, corresponding to the intra-tile selected by the NLP, is implemented as an independent, fully unrolled function with no communication with off-chip memory. Data for these tasks resides on-chip. If the task includes reduction loops, inter-tile reduction loops are incorporated and pipelined within the function. The pipeline’s II is strictly greater than 1 due to reduction loop dependencies, but other operations within the statements are effectively pipelined. Padding is managed at the intra-tile level. For non-reduction loops, padding does not alter the intra-task, allowing computation of padding values without unnecessary resource usage for minor latency improvements. In contrast, for reduction loops, full tile computations are performed first, followed by adjusting the intra-tile loop to handle padding and compute partial tiles accurately.
Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
Concurrent Execution  With pragma dataflow, independent tasks run concurrently without rewriting. Computation begins as soon as the receiving task’s shared buffer has enough data.
SLR  The NLP determines the SLR on which each task will be executed. A separate C++ file is generated for each SLR, and data transfers between SLRs are handled using ap_axiu.
VI Evaluation
VI-A Setup and Experimental Evaluation
We evaluated our method using kernels from Polybench/C 4.2.1 [56] with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow [4]. This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we comare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE. We execute the designs for these frameworks using identical parameters as outlined in Section VI-A. For ScaleHLS [79] and Allo [14], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory. Their kernels assume data are already in on-chip memory, so we modified their code to include off-chip to on-chip data transfers. Allo’s 2mm and 3mm kernels already handle this transfer, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE [1]. Sisyphus generated designs are generated using parameters consistent with the paper [60], and ScaleHLS designs are compiled using their open-source compiler [2]. For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
VI-B Comparison
The average time to solve the NLP and generate the designs presented in this section is 31.7 seconds, with a maximum of 178.416 seconds for 3mm with 3 SLRs.
RTL Simulation  Table II presents the RTL simulation results for Prometheus, Sisyphus [60], AutoDSE [67], ScaleHLS [79] (S-HLS), and Allo [14]. The last lines show the average and geometric mean performance improvement (PI) of Prometheus across evaluated kernels. Results marked with * are from the Vitis HLS report, showing minimum latency as an optimistic estimate. The RTL simulation for 3mm with Allo was incomplete after two days.
Prometheus consistently achieves superior QoR across all evaluated kernels compared to other frameworks. While Sisyphus [60], designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm. This disparity is due to Sisyphus lacking concurrent execution capabilities for independent matrix multiplications. The performance advantage of Prometheus stems from both concurrent task execution and efficient overlapping of computation and communication.
Kernel Ours Sisyphus S-HLS Allo AutoDSE
2mm 308.38 195.09 37.13 46.58 0.41*
3mm 368.36 178.97 43.04 60.40* 1.74*
Atax 3.56 2.32 1.58 1.96 1.97*
Bicg 15.41 2.32 1.70 14.17 0.99*
Gemm 419.14 227.09 40.53 37.50 110.81*
Gesummv 10.21 2.28 1.78 8.85 1.98*
Mvt 14.65 5.54 7.39 8.77 7.80*
Symm 212.20 200.30 0.06 16.72 14.68*
Syr2k 133.77 75.01 0.04 20.59 12.35*
Syrk 316.25 211.00 0.54 41.11 23.16*
Trmm 193.47 166.72 0.07 5.10 0.02*
PI (Avg) 2.39x 927.20x 8.58x 973.14
PI (gmean) 2.03x 48.03x 4.92x 25.82
TABLE II: Comparison of the throughput (GF/s) of various optimization framework with RTL Simulation
Although for bicg, Allo and Prometheus do not use the same code transformation, the results are similar. Allo retains the original code structure by permuting the reduction loop outermost. The non-reduction loop is fully unrolled, and the reduction loop is pipelined. Prometheus partially unrolls both loops and pipelines the reduction loop.
On Board Evaluation  Table III presents the results for the on-board evaluation. The column $T(ms)$ indicates the kernel execution time in milliseconds, GF/s represents the throughput in Giga Floating Operations per second, and the resource usage is detailed in the thousands for both LUTs (L) and FFs. URAM utilization is excluded as no kernels use it. Column F (MHz) shows the frequency achieved by each design, with a target frequency of 220 MHz for all designs.
For Bicg and Atax, targeting 60% of resources on one SLR caused congestion, requiring regeneration with a 55% constraint. The 3mm bitstream from AutoDSE succeeded only with a 15% constraint.
Kernel T (ms) GF/s DSP BRAM L(K) FF(K) F (MHz)
1 SLR Sisyphus 2mm 1.20 30.57 556 510 213 276 220
3mm 1.52 29.89 984 611 230 300 220
Atax 0.62 1.03 173 450 240 250 220
Bicg 0.63 1.02 173 451 238 265 217
1 SLR AutoDSE 2mm 92.25 0.40 963 353.5 287 292 205
3mm 110.34 0.41 1117 470 278 306 220
Atax 2.88 0.22 452 630.5 170 212 220
Bicg 1.13 0.56 196 867.5 168 217 214
1 SLR Ours 2mm 0.56 65.13 1941 635.5 371 454 216
3mm 0.87 51.95 1551 635.5 342 423 220
Atax 0.24 2.62 1081 533.5 234 287 184
Bicg 0.15 4.04 732 311.5 250 302 220
3 SLR Ours 2mm 0.29 125.54 2752 546 428 549 220
3mm 0.34 134.07 4379 600 684 840 207
Atax 0.20 3.10 1823 634.5 405 539 137
Bicg 0.14 4.34 1226 241 291 380 177
TABLE III: On-board evaluation comparison
Prometheus achieves a remarkable 77.16x performance improvement over AutoDSE. This gain is accompanied by an average increase in resource utilization: 2.38x more DSPs, 1.08x more BRAM, 1.36x more LUTs, and 1.42x more FFs, reflecting the trade-offs made to achieve such high efficiency. When compared to Sisyphus, Prometheus demonstrates a notable 2.59x performance boost, leveraging an average of 3.88x more DSPs, 1.04x more BRAM, 1.31x more LUTs, and 1.33x more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
Table IV shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 [56] code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task is precise.
Permutations are evident in the implementations of 3mm, Atax, and Bicg. For 2mm and 3mm, the NLP opted to fully transfer array B instead of overlapping computation and load, as the load operation had higher latency than computation. In 2mm, the NLP retains the original loop order. Array tmp is transferred from the first task to the second, with both tasks iterating over the first dimension using loop $i$. This enables a $4\times 32$ data tile to be sent to the second task, which starts computation once a $4\times 192$ tile of tmp(FT1) is filled.
Fused Statement Loop   Order Data Tile Size
2mm FT0: S0, S1, FT1: S2, S3 FT0: i,j,k, FT1: i,j,k tmp(FT0): $4\times 32$, B: $212\times 192$, A: $4\times 212$, D: $4\times 32$, C:$192\times 32$, tmp(FT1): $4\times 192$
3mm FT0: S0, S1, FT1: S2, S3, FT2: S4, S5 FT0: i,j,k, FT1: j,i,k FT2: j,i,k E(FT0): $9\times 16$, A: $9\times 200$, B: $200\times 192$, F(FT1): $10\times 10$, C: $10\times 224$, D: $224\times 10$, G: $9\times 10$, E(FT2): $9\times 192$, F(FT2): $192\times 10$,
Atax FT0: S1, S2, FT1: S0, S3 FT0: i,j, FT1: j,i tmp(FT0): $56$, y: $16$, A(FT0): $392\times 416$, A(FT1): $392\times 16$, tmp(FT1): $392$
Bicg FT0: S1, S2, FT1: S0, S3 FT0: j,i FT1: i,j s: $16$, A (FT0): $416\times 16$, r: $416$, q: $10$, A (FT1): $10\times 400$, p: $400$,
TABLE IV: Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table III for 1 SLR
As other frameworks cannot generate bitstreams for 3 SLRs without human intervention, we compare results for 1 and 3 SLRs using our framework. For 2mm and 3mm, performance improves due to increased resource utilization. However, for atax and bicg, where the bottleneck is memory transfer between off-chip and on-chip rather than parallelism, the improvement is negligible.
VII Related Work
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks [34], Dennis Dataflow [22], synchronous dataflow languages [42, 11], and for FPGA applications [53, 12, 3, 25, 10, 5, 75, 14]. DaCe [10] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Flower [5] automates FPGA dataflow development but limits parallelism. Frameworks like [14, 75], built on HeteroCL [38], optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays [71, 20, 33, 39, 9] offer efficient computation for specific patterns but lack generalization. Application-specific frameworks [16, 29, 51, 21, 63, 23, 36, 6] demonstrate dataflow advantages but do not generalize across domains.
Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion [58, 59, 60], bottleneck DSE [67], and GNN-based latency and resource estimation [66, 64, 65, 8, 68, 73, 74]. However, these approaches lack effective integration of dataflow optimization techniques.
Code Transformation  Code transformation has been explored for CPUs [13, 55, 37, 7], GPUs [70], and FPGAs [57, 84, 83, 44, 48, 47, 46, 17]. Pluto [13] minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations [57] leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works [84, 83] selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While [44, 48, 47, 46, 17] focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The [79, 38, 31, 14, 78] compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper [60] described in Section II generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
Tiling and Padding  Tiling is essential for balancing computation and communication. While prior works [49, 45] use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling [45] focus on CPUs, while Wedler [61] optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses [30, 52] and improving memory transfers [72], but its use for varying unroll factors on FPGAs remains underexplored.
VIII Conclusion
In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework. By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to maximize throughput within stringent resource constraints. Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.