### v2

# Holistic Optimization Framework for FPGA Accelerators

Abstract
Customized accelerators have transformed modern computing by delivering substantial gains in energy efficiency and performance through hardware specialization. Field-Programmable Gate Arrays play a critical role in this paradigm, offering unmatched flexibility and high-performance potential. High-Level Synthesis and source-to-source compilers have simplified Field-Programmable Gate Arrays design by translating high-level programming languages into hardware descriptions enriched with directives. However, achieving high Quality of Results remains a significant challenge, requiring intricate code transformations, strategic directive placement, and optimized data communication.
This paper presents Prometheus, a comprehensive framework that integrates key optimizations, including task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution, into a unified design space. By leveraging Non-Linear Problem methodologies, Prometheus explores this design space to deliver efficient solutions under strict resource constraints, facilitating automatic bitstream generation. The framework addresses limitations in existing methods and demonstrates superior performance, achieving high Quality of Results across complex computation-bound applications.
I Introduction
The rise of customized accelerators has transformed modern computing, delivering substantial improvements in energy efficiency and performance through hardware specialization. Field-Programmable Gate Arrays (FPGAs) have emerged as a particularly versatile solution, valued for their flexibility and ability to achieve high performance across diverse applications. High-level synthesis (HLS) tools and source-to-source compilers [20, 80, 75, 32, 61, 49, 76, 31, 78, 37, 79, 42, 59, 9] have simplified hardware design by translating high-level programming languages (e.g., C++, Python) enriched with hardware-specific directives into synthesizable hardware descriptions. However, the effectiveness of these tools depends heavily on the quality of the input code and the strategic application of hardware directives, making it challenging to achieve high Quality of Results (QoR) without significant manual effort.
Despite advancements in the field, designing optimized FPGA solutions remains a complex task. Existing approaches [57, 58, 66, 65, 63, 25, 81, 65, 37, 78, 77, 59, 70, 9, 31] often address specific aspects of FPGA design in isolation, leaving gaps in areas such as efficient code transformation, effective concurrency management, and resource optimization. Moreover, challenges such as overlapping computation and communication, balancing intra-task parallelism, and managing FPGA resource constraints—including multi-Super Logic Region (SLR) architectures—further complicate the development process. These challenges are exacerbated by the lack of holistic frameworks that integrate multiple optimization strategies while ensuring synthesizable output.
Two commonly used paradigms—shared buffer and dataflow—have shown promise in improving FPGA performance. Shared buffers enable data reuse and promote parallelism by caching information on-chip, which is particularly beneficial for computation-bound kernels. However, their efficacy is often constrained by the limited on-chip memory, preventing full data caching and hindering the ability to overlap computation and communication effectively. On the other hand, the dataflow paradigm facilitates efficient data movement and task concurrency by leveraging FIFO (First-In-First-Out) communication. While this approach is effective for memory-bound kernels, it struggles to exploit shared memory and parallelism within tasks, and its independent task compilation can limit resource sharing opportunities.
This work addresses these challenges by focusing on optimizing HLS for FPGA design, with a particular emphasis on reducing latency while adhering to resource constraints. The goal is to transform affine high-level code (e.g., Listing 1) into an SLR-aware HLS-C++ design that achieves high QoR and facilitates automated bitstream generation. Achieving this requires addressing multiple challenges simultaneously, including task-to-SLR mapping, balancing computation and communication, and efficiently regenerating code to adapt to resource conflicts or spatial congestion. Existing methods fall short of providing an integrated solution capable of achieving these objectives, highlighting the need for a unified approach.
To address these limitations, we propose Prometheus, an integrated framework that combines the benefits of shared buffer and dataflow paradigms to optimize HLS-based FPGA design. The key contributions of this work include:
  • A unified design space supporting tiled dataflow architectures, task fusion, tiling strategies, loop permutation, overlapping computation and communication, and concurrent task execution.
  • The application of a Non-Linear Problem (NLP) formulation to discover optimal design parameters within strict resource constraints, enabling efficient bitstream generation.
  • An end-to-end automated framework that generates OpenCL host code for correctness verification, performs Vitis HLS simulations, and produces the final bitstream.
  • Rigorous evaluation through RTL simulation and on-board evaluation, demonstrating superior QoR compared to state-of-the-art tools like AutoDSE [66], Sisyphus [59], and ScaleHLS [78].
II Background and Motivation
II-A HLS Optimization
HLS optimizations rely heavily on directives such as unroll, pipeline, and array_partitioning, which play a pivotal role in enhancing performance. The unroll pragma duplicates the loop body in the RTL design to enable parallel execution of loop iterations, while pipeline overlaps loop operations, reducing latency based on the initiation interval (II). Similarly, array_partitioning divides arrays into on-chip memory banks, allowing simultaneous access and facilitating more effective unrolling.
Advanced optimizations also leverage shared buffers to reduce off-chip memory transfers, ping-pong buffering to overlap data transfer and computation, and coarse-grained parallelism through duplicated processing elements [19]. Modern AMD/Xilinx FPGAs support data transfers with bit widths up to 512 bits, determined by problem dimensions. For instance, transferring 216 floats using a 256-bit width (8 floats per transfer) requires 27 transfers.
HLS tools like Vitis HLS support the dataflow pragma, which organizes computations into independent actors communicating through FIFO queues [40]. This structure enables concurrent execution by allowing each actor to process inputs and generate outputs as soon as data becomes available. Performance evaluations often combine estimation (e.g., Vitis HLS reports), RTL simulation, and on-board evaluation. While estimations and simulations avoid place-and-route overheads, on-board evaluations impose stricter constraints for generating deployable bitstreams.
II-B Limitations of Existing Methods
The limitations of current HLS optimization frameworks are illustrated through the 3mm kernel, which performs two independent matrix multiplications followed by a dependent multiplication. As shown in Listing 1, each matrix multiplication involves nested loops that can benefit from parallelism, unrolling, and communication-computation overlap.
1for (i = 0; i < 180; i++) // MM 1
2 for (j = 0; j < 190; j++) {
3 E[i][j] = 0.0; // S0
4 for (k = 0; k < 200; ++k)
5 E[i][j] += A[i][k] * B[k][j];} // S1
6for (i = 0; i < 190; i++) // MM 2
7 for (j = 0; j < 210; j++) {
8 F[i][j] = 0.0; // S2
9 for (k = 0; k < 220; ++k)
10 F[i][j] += C[i][k] * D[k][j];} // S3
11for (i = 0; i < 180; i++) // MM 3
12 for (j = 0; j < 210; j++) {
13 G[i][j] = 0.0; // S4
14 for (k = 0; k < 190; ++k)
15 G[i][j] += E[i][k] * F[k][j];} // S5
Listing 1: Code of 3mm
Shared Buffer  Frameworks such as AutoDSE [66], Sisyphus [59], ScaleHLS [78], and others [65, 63, 58, 57] optimize 3mm primarily through sequential execution of matrix multiplications. This approach enables reuse of DSP resources but fails to leverage opportunities for concurrent execution, such as running the first two multiplications in parallel. Current frameworks transfer arrays at different loop levels to manage on-chip memory but lack automatic mechanisms to overlap computation with communication effectively.
Although shared buffering could support concurrent execution of independent tasks, its application remains limited. Extending concurrency to all three multiplications would require task fusion, which may constrain compiler optimizations.
Dataflow  The dataflow paradigm, as implemented in frameworks like StreamHLS [9], divides computations into independent tasks, with communication managed through FIFOs. For the 3mm kernel in Listing 1, dataflow enables concurrent execution of the first two matrix multiplications, while the third begins as soon as its inputs are ready. This approach overlaps computation and communication, effectively managing task dependencies. As 3mm is computation-bound, increasing parallelism is essential. However, parallelism in pure dataflow methods is constrained by the FIFO bit width of the receiving task and the number of FIFOs. Increasing FIFO bit widths or adding more FIFOs to handle higher concurrency often results in routing congestion, making it challenging to generate a valid bitstream. Additionally, the lack of shared memory prevents data reuse across tasks, leading to repeated off-chip transfers or redundant buffers.
Mixed Methods  Mixed approaches, such as those used by Allo [15] and HIDA [77], combine shared buffering and dataflow paradigms. Allo extends HeteroCL [37] within the MLIR framework [39], supporting code transformation and dataflow scheduling. However, these methods lack automatic data-tile selection and pragma insertion, limiting their ability to optimize QoR effectively.
General Limitations  Most existing HLS frameworks fail to address key challenges such as automatic ping-pong buffering, tile size adjustments, and padding to maximize bit widths for communication efficiency. Unroll factors are often restricted to evenly divide problem dimensions, which can lead to inefficiencies with partial tiles. Expanding these factors and incorporating dynamic adjustments could significantly enhance parallelism and performance.
II-C Overview of Prometheus
We present Prometheus, a unified optimization framework designed to efficiently explore design spaces using a Non-Linear Problem (NLP) cost model. Prometheus simplifies the development of FPGA designs by offering an engineer-friendly interface with adjustable parameters such as tile size, array partitioning, and unroll factors. It is SLR-aware, ensuring effective task distribution across Super Logic Regions (SLRs), which simplifies bitstream generation while optimizing the balance between performance and resource utilization.
The pseudo-code in Listing 2 demonstrates how Prometheus processes the 3mm kernel. Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 1. An example of such a task is illustrated in Listing 3.
Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 1), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
1/*************** Fused Task 0 ****************/
2float B[204][192]; load_B(B);
3for (i0 = 0; i0 < 18; i0++)//inter-tile loop
4 float A[10][204]; load_A(A);
5 for (j0 = 0; j0 < 6; j0++){//inter-tile loop
6 float E[10][32];
7 task0(E); // S0 + intra-tile loops
8 for (k0 = 0; k0 < 51; ++k0)//inter-tile loop
9#pragma HLS pipeline II=3
10 task1(E, A, B); // S1 + intra-tile loops
11 store_E(E); sent_E(E);
12/*************** Fused Task 1 ****************/
13float C[190][222]; load_C(C);
14for (j0 = 0; j0 < 7; j0++)//inter-tile loop
15 float D[222][32]; load_D(D);
16 for (i0 = 0; i0 < 10; i0++){//inter-tile loop
17 float F[19][32];
18 task2(F); // S2 + intra-tile loops
19 for (k0 = 0; k0 < 74; ++k0)//inter-tile loop
20#pragma HLS pipeline II=3
21 task3(F, C, D); // S3 + intra-tile loops
22 store_F(F); sent_F(F);
23/*************** Fused Task 2 ****************/
24float E[180][192];
25for (j0 = 0; j0 < 7; j0++)//inter-tile loop
26 float F[192][32];
27 receive_F(F);
28 for (i0 = 0; i0 < 18; i0++){//inter-tile loop
29 float G[10][32];
30 receive_E(E);
31 task4(G); // S4 + intra-tile loops
32 for (k0 = 0; k0 < 32; ++k0)//inter-tile loop
33#pragma HLS pipeline II=3
34 task5(G, E, F); // S5 + intra-tile loops
35 store_G(G);
Listing 2: Pseudo Code of 3mm generated by Prometheus
Efficient computation-communication overlap is achieved through ping-pong buffering, while concurrent task execution and FIFO-triggered dependent tasks maximize overall performance. Fused Tasks 0 and 1 execute concurrently, while Fused Task 2 begins as soon as the data tiles for $F$ and $E$ become available.
For the 3mm kernel optimized with RTL simulation, Prometheus achieved a throughput of 368.36 GF/s, significantly surpassing state-of-the-art frameworks such as Sisyphus (178.97 GF/s), Allo (60.40 GF/s), ScaleHLS (43.04 GF/s), and AutoDSE (1.74 GF/s). These results demonstrate the effectiveness of Prometheus in overcoming the limitations of existing methods and achieving superior QoR.
III Code Transformation
In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time. This compile-time awareness enables us to construct a precise model that facilitates rigorous optimizations. To leverage FPGA parallelism, we implement an acyclic dataflow graph, ensuring parent nodes do not receive data from their children. While this constraint limits graph configurations and may increase resource usage, it reduces overall latency. To support this structure, we inline [16] each function in the input code to generate the required acyclic graph. Our primary objective is to minimize latency within resource constraints by overlapping communication and computation within tasks and executing independent tasks concurrently. To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section IV.
Dependency Graph Creation  Our process starts with affine C/C++ code as input, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC [68] verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, we construct a dependency graph using PoCC [53]. PoCC provides the necessary information about the schedule and dependencies, enabling us to build the graph. In this graph, the nodes represent tasks, while the edges capture data communication arising from inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. This ensures that each tile’s output is handled (loaded, computed, and either stored or transmitted) only once. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
Prometheus’s solution space includes:
Data-tiling and Padding  In the fused dependency graph, edges represent data communication between tasks as well as between tasks and off-chip memory, involving the transfer of data tiles with specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. For each array and dimension iterated by the loop $l$, the tile factor is a common choice that influences all arrays iterated by this loop.
To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. In Listing 2, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
Fine-grained Parallelism  Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops. Using ISCC [68], we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is not feasible, the inter-tile loop retains its legal position. For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction. For instance, in Listing 2, the inter-tile loops on lines 3 and 5 iterate over task0 and task1. The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 2, additions take 3 cycles, resulting in $II=3$.
Loop Order  Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP (cf. Section IV). As shown in Listing 2, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 1.
Automatic Overlapping of Communication and Computation  To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section IV). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only or write-only arrays, and triple-buffering for arrays that are both read and written. Following [19], we perform an initial load, then overlap loading the next tile with computing the current one.
Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks  Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution. This concurrency significantly increases overall performance.
Memory Transfer  Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory transfer.
IV NLP Formulation
To determine the tile size, loop order, bit width, and memory transfers, we formulate a cost model as a NLP problem aimed at minimizing overall latency. This approach builds upon the methodology proposed by Pouget et al. [58, 57, 59], which we have adapted to address our specific requirements and constraints. In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section III. We employ PoCC [53] to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement. ISCC [68] then generates all legal permutations for each loop body.
Table I delineates the sets, variables, and constants utilized in our NLP formulation.
Constants Description
$II_{l}$ II of the loop $l$
$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
$TC^{l}_{ori}$ Original Trip Count of the loop $l$
$f_{a,l}$ Footprint of the array $a$ if transferred to on-chip after the loop $l$
$N_{FT}$ Number of fused task
$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation $op$
$DSP$ Number of DSP available for the FPGA used
$max_{part}$ Maximum array partitioning
$SLR$ Number of SLR available for the FPGA used
Variables Description
$TC^{l}_{intra}$, $TC^{l}_{inter}$ TC of the loop $l$ for the intra and inter tile
$TC^{l}$ Trip Count of the loop $l$ after padding
$S^{last}_{a}$ Size of the last dimension of the array $a$ transferred on-chip
$BW_{a}$ Bit width of the array $a$
$t_{a,l},d_{a,l}$ Boolean to know if the array $a$ is transferred and defined (respectively) on-chip after the loop $l$ in the inter-tile
$p_{i}^{l}$ Position of the loop $l$ under the $i$-th permutation
$slr_{t}$ ID of the SLR use by the task $t$
Sets Description
$\mathcal{L},\mathcal{A},\mathcal{S}$ The set of loops, arrays and statements
$\mathcal{L}_{s}$ The set of loops which iterate the statement $s$
$\mathcal{L}_{s}^{red}$ The set of reduction loops which iterate the statement $s$
$\mathcal{L}_{a}^{last}$ The set of loop which iterate the last dimension of the array $a$
$\mathcal{L}_{inter},$ $\mathcal{L}_{intra}$ The set of loops which belong to the inter-tile and intra-tile respectively
$\mathcal{B}$ Set of possible burst size for the data type
$\mathcal{C}_{a_{d}}$ The set of loops which iterates the array $a$ at the dimension $d$
$AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
$\mathcal{P}_{s}$ All permutation of the loops which iterate the statement $s$
$\mathcal{F}_{i}$ The set of statements which belong to the fused task $i$
$\mathcal{T}$ The set of tasks in the dataflow
TABLE I: Overview of the constant, variable and set employed in formulating the NLP
IV-A Constraints
We now describe the constraints by using the code of Listings 1, 2 and 3.
1for (int j1 = 0; j1 < 32; j1++) {
2#pragma HLS unroll
3 for (int i1 = 0; i1 < 19; i1++) {
4#pragma HLS unroll
5 for (int k1 = 0; k1 < 3; k1++) {
6#pragma HLS unroll
7 j=j0*32+j1; i=i0*19+i1; k=k0*3+k1;
8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 3: Pseudo Code of the task_3
Data-tiling and Unroll Factor  The intra-tile transformation, as explained in Section III, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 3, for the array $F$, the loop $j$ (line 7 in Listing 1) has been split into $j0$ (line 14 in Listing 2) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
$\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$ (1)
$(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$ (2)
Bit Width  $\mathcal{B}$ denotes the number of elements that can be transferred simultaneously, determined by the bit width and the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
$\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$ (3)
Permutation  Equation 4 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
$\forall i\in\llbracket 0,N_{FT}\llbracket,\forall(ft_{0},ft_{1})\in\mathcal{F}_{i}^{2},$
$\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$ (4)
Transfer and Reuse  Equation 5 permits the selection of a single level where each array can be defined (and reused) and transferred. Equation 6 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
The array $E$, defined on line 24 in Listing 2, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 6 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 2, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
$\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$ (5)
$\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$ (6)
On-chip Memory  Equation 7 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
$\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$ (7)
with $N_{a}$ being the number of double buffer for the array $a$.
Array Partitioning  Equation 8 limits the maximum partitioning of each array. This partitioning is crucial as it impacts the maximum unroll factor, necessitating the distribution of data across different BRAM banks under fully unrolled loops, thereby influencing the utilization of BRAMs. Equation 9 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
Array D in Listing 3 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
$\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$ (8)
$\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$ (9)
DSP Utilization  Equation 10 constrains the number of DSPs used based on the available DSP resources. In contrast to [58, 57, 59], we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously. Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
$\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$ (10)
SLR Selection  For each task, the NLP determines the SLR on which the task will be implemented (Equation 11). Additionally, Equations 7 and 10 are applied to each SLR to manage resource allocation per SLR effectively.
$\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$ (11)
IV-B Objective Function
The objective function models the dependency graph of the fused task. Independent tasks execute simultaneously, with latency set by the longest task. Dependent tasks overlap partially, with the second task starting once it receives enough data from the first. The latency is the maximum of both tasks’ durations plus the “shift”, the interval before the second task starts. For our 3mm example, the graph is shown in Figure 1, and the corresponding latency is:
$\left.\begin{cases}&Lat(T_{0},T_{1})=\max(Lat(T_{0}),Lat(T_{1})+shift_{T_{0},T_{1}})\\ &Lat(T_{2},T_{3})=\max(Lat(T_{2}),Lat(T_{3})+shift_{T_{2},T_{3}})\\ &Lat(T_{0},..,T_{4})=\max(Lat(T_{0},T_{1}),Lat(T_{2},T_{3}),Lat(T_{4}))\\ &Lat=\max(Lat(T_{0},..,T_{4}),Lat(T_{5})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\max(shift_{T_{1},T_{5}},shift_{T_{3},T_{5}},shift_{T_{4},T_{5}}))\\ \end{cases}\right.$
T0T1T2T3T4T5
Figure 1: Dataflow graph of 3mm maximally distributed
The latency of the intra-task for a statement $s$ is determine similarly to [58, 57, 59] with Equation 12. And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 13.
$Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$ (12)
$Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$ (13)
For each level $n\geq 0$ (with the outermost level being 0) of the inter-tile, we consider the shifting of load, compute, and store operations to enable overlapping of computation and communication. The overlap factor $\alpha$ is set to 1 for read-only operations and 2 for read and store operations:
$Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$ (14)
And then, the final objective function is computed using the dataflow graph, where the latency of the task is established by the latency of the outermost level, denoted as $Lat_{0}$.
V Code Generation
Prometheus takes as input affine C/C++ code and automatically produces an HLS-C++ file, OpenCL host files, and all necessary files for code verification, RTL simulation, and bitstream generation. The NLP described in Section IV gives the parameters of the space such as loop order, tiling factor, etc. However, in order to be efficient, the code generation needs some specific rules.
Communication with Off-Chip Memory  To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler. Furthermore, to guarantee that the load operation transfers data with the correct bit width, we automatically restructure the data in off-chip memory to enable sequential loading. Once data accumulates in the FIFO, a read function moves it to the shared data-tile buffer, whose size is determined by the NLP. For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
Communication between the Fused Task  The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
Intra-Task  Each intra-task, corresponding to an intra-tile selected by the NLP, is implemented as a fully unrolled, independent function without communication with off-chip memory. Data for these tasks resides entirely on-chip.
If the task involves reduction loops, inter-tile reduction loops are integrated into the function and pipelined. While the pipeline initiation interval (II) is greater than 1 due to reduction dependencies, other operations in the statement are pipelined efficiently.
Padding is handled at the intra-tile level. Non-reduction loops remain unchanged, allowing computation of padding values without excessive resource usage. For reduction loops, full tiles are computed first, and the intra-tile loop is adjusted to handle padding for partial tiles accurately.
Inter-Tile Loop  For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads or writes are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
Concurrent Execution  Using the dataflow pragma, independent tasks execute concurrently without requiring manual rewrites. Computation in a receiving task begins as soon as its shared buffer contains sufficient data.
SLR Management  The NLP determines the SLR ID for each task, specifying where it will be executed. Prometheus generates a separate C++ file for each SLR, and data transfers between SLRs are managed via ap_axiu streams, ensuring efficient communication and minimizing transfer overhead.
VI Evaluation
VI-A Setup and Experimental Evaluation
We evaluated our method using kernels from Polybench/C 4.2.1 [55] with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow [4]. This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we compare our work with Allo, ScaleHLS, Sisyphus, and AutoDSE. We do not include a comparison with Stream-HLS [9] due to the unavailability of its source code. For ScaleHLS [78] and Allo [15], we apply the same method as our framework by padding memory transfers to attain the maximal bit width (512 bits) from off-chip to on-chip memory. Their kernels assume data are already in on-chip memory, so we modified their code to include off-chip to on-chip data transfers. Allo’s 2mm and 3mm kernels already handle this transfer, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE [1]. Sisyphus generated designs are generated using parameters consistent with the paper [59], and ScaleHLS designs are compiled using their open-source compiler [2]. For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
VI-B Comparison
The average time to solve the NLP and generate the designs presented in this section is 31.70 seconds, with a maximum of 178.41 seconds for 3mm with 3 SLRs.
RTL Simulation  Table II presents the RTL simulation results for Prometheus, Sisyphus [59], AutoDSE [66], ScaleHLS [78] (S-HLS), and Allo [15]. The last lines show the average and geometric mean performance improvement (PI) of Prometheus across evaluated kernels. Results marked with * are from the Vitis HLS report, showing minimum latency as an optimistic estimate. The RTL simulation for 3mm with Allo was incomplete after two days.
Prometheus consistently achieves superior QoR across all evaluated kernels compared to other frameworks. While Sisyphus [59], designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm. This disparity is due to Sisyphus lacking concurrent execution capabilities for independent matrix multiplications. The performance advantage of Prometheus stems from both concurrent task execution and efficient overlapping of computation and communication.
Kernel Ours Sisyphus S-HLS Allo AutoDSE
2mm 308.38 195.09 37.13 46.58 0.41*
3mm 368.36 178.97 43.04 60.40* 1.74*
Atax 3.56 2.32 1.58 1.96 1.97*
Bicg 15.41 2.32 1.70 14.17 0.99*
Gemm 419.14 227.09 40.53 37.50 110.81*
Gesummv 10.21 2.28 1.78 8.85 1.98*
Mvt 14.65 5.54 7.39 8.77 7.80*
Symm 212.20 200.30 0.06 16.72 14.68*
Syr2k 133.77 75.01 0.04 20.59 12.35*
Syrk 316.25 211.00 0.54 41.11 23.16*
Trmm 193.47 166.72 0.07 5.10 0.02*
PI (Avg) 2.39x 927.20x 8.58x 973.14
PI (gmean) 2.03x 48.03x 4.92x 25.82
TABLE II: Comparison of the throughput (GF/s) with RTL Simulation
Although for bicg, Allo and Prometheus do not use the same code transformation, the results are similar. Allo retains the original code structure by permuting the reduction loop outermost. The non-reduction loop is fully unrolled, and the reduction loop is pipelined. Prometheus partially unrolls both loops and pipelines the reduction loop.
On Board Evaluation  Table III presents the results for the on-board evaluation. The column $T(ms)$ indicates the kernel execution time in milliseconds, GF/s represents the throughput in Giga Floating Operations per second, and the resource usage is detailed in the thousands for both LUTs (L) and FFs. URAM utilization is excluded as no kernels use it. Column F (MHz) shows the frequency achieved by each design, with a target frequency of 220 MHz for all designs.
For Bicg and Atax, targeting 60% of resources on one SLR caused congestion, requiring regeneration with a 55% constraint. The 3mm bitstream from AutoDSE succeeded only with a 15% constraint.
Kernel T (ms) GF/s DSP BRAM L(K) FF(K) F (MHz)
1 SLR Sisyphus 2mm 1.20 30.57 556 510 213 276 220
3mm 1.52 29.89 984 611 230 300 220
Atax 0.62 1.03 173 450 240 250 220
Bicg 0.63 1.02 173 451 238 265 217
1 SLR AutoDSE 2mm 92.25 0.40 963 353.5 287 292 205
3mm 110.34 0.41 1117 470 278 306 220
Atax 2.88 0.22 452 630.5 170 212 220
Bicg 1.13 0.56 196 867.5 168 217 214
1 SLR Ours 2mm 0.56 65.13 1941 635.5 371 454 216
3mm 0.87 51.95 1551 635.5 342 423 220
Atax 0.24 2.62 1081 533.5 234 287 184
Bicg 0.15 4.04 732 311.5 250 302 220
3 SLR Ours 2mm 0.29 125.54 2752 546 428 549 220
3mm 0.34 134.07 4379 600 684 840 207
Atax 0.20 3.10 1823 634.5 405 539 137
Bicg 0.14 4.34 1226 241 291 380 177
TABLE III: On-board evaluation comparison
Prometheus achieves a remarkable 77.16x performance improvement over AutoDSE. This gain is accompanied by an average increase in resource utilization: 2.38x more DSPs, 1.08x more BRAM, 1.36x more LUTs, and 1.42x more FFs, reflecting the trade-offs made to achieve such high efficiency. When compared to Sisyphus, Prometheus demonstrates a notable 2.59x performance boost, leveraging an average of 3.88x more DSPs, 1.04x more BRAM, 1.31x more LUTs, and 1.33x more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
Table IV shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 [55] code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task (defined in the second column) is precise.
Permutations are evident in the implementations of 3mm, Atax, and Bicg. For 2mm and 3mm, the NLP opted to fully transfer array B instead of overlapping computation and load, as the load operation had higher latency than computation. In 2mm, the NLP retains the original loop order. Array tmp is transferred from the first task to the second, with both tasks iterating over the first dimension using loop $i$. This enables a $4\times 32$ data tile to be sent to the second task, which starts computation once a $4\times 192$ tile of tmp(FT1) is filled.
Fused Statement Loop   Order Data Tile Size
2mm FT0: S0, S1, FT1: S2, S3 FT0: i,j,k, FT1: i,j,k tmp(FT0): $4\times 32$, B: $212\times 192$, A: $4\times 212$, D: $4\times 32$, C:$192\times 32$, tmp(FT1): $4\times 192$
3mm FT0: S0, S1, FT1: S2, S3, FT2: S4, S5 FT0: i,j,k, FT1: j,i,k FT2: j,i,k E(FT0): $9\times 16$, A: $9\times 200$, B: $200\times 192$, F(FT1): $10\times 10$, C: $10\times 224$, D: $224\times 10$, G: $9\times 10$, E(FT2): $9\times 192$, F(FT2): $192\times 10$,
Atax FT0: S1, S2, FT1: S0, S3 FT0: i,j, FT1: j,i tmp(FT0): $56$, y: $16$, A(FT0): $392\times 416$, A(FT1): $392\times 16$, tmp(FT1): $392$
Bicg FT0: S1, S2, FT1: S0, S3 FT0: j,i FT1: i,j s: $16$, A (FT0): $416\times 16$, r: $416$, q: $10$, A (FT1): $10\times 400$, p: $400$,
TABLE IV: Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table III for 1 SLR
As other frameworks cannot generate bitstreams for 3 SLRs without human intervention, we compare results for 1 and 3 SLRs using our framework. For 2mm and 3mm, performance improves due to increased resource utilization. However, for atax and bicg, where the bottleneck is memory transfer between off-chip and on-chip rather than parallelism, the improvement is negligible.
VII Related Work
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks [34], Dennis Dataflow [23], synchronous dataflow languages [41, 12], and for FPGA applications [52, 13, 3, 26, 11, 5, 74, 15]. DaCe [11] introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Stream-HLS [9] automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory. Additionally, it offers very limited opportunities for parallelism. Flower [5] automates FPGA dataflow development but limits parallelism. Frameworks like [15, 74], built on HeteroCL [37], optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays [70, 21, 33, 38, 10] offer efficient computation for specific patterns but lack generalization. Application-specific frameworks [17, 29, 50, 22, 62, 24, 35, 6] demonstrate dataflow advantages but do not generalize across domains. RapidStream-TAPA [27, 28] enhances the performance of dataflow designs and automates SLR placement. However, it requires an optimized kernel with a dataflow structure as input.
Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion [57, 58, 59], bottleneck DSE [66], and GNN-based latency and resource estimation [65, 63, 64, 8, 67, 72, 73]. However, these approaches lack effective integration of dataflow optimization techniques.
Code Transformation  Code transformation has been explored for CPUs [14, 54, 36, 7], GPUs [69], and FPGAs [56, 83, 82, 43, 47, 46, 45, 18]. Pluto [14] minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations [56] leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works [83, 82] selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While [43, 47, 46, 45, 18] focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The [78, 37, 31, 15, 77] compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper [59] described in Section II generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
Tiling and Padding  Tiling is essential for balancing computation and communication. While prior works [48, 44] use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling [44] focus on CPUs, while Wedler [60] optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses [30, 51] and improving memory transfers [71], but its use for varying unroll factors on FPGAs remains underexplored.
VIII Conclusion
In this work, we introduce Prometheus, a comprehensive FPGA performance enhancement framework. By integrating techniques like task fusion, automatic tiling, loop permutation, and overlap of communication and computation, Prometheus is intended to minimize latency within stringent resource constraints. Our methodology utilizes NLP to navigate complex design spaces and automate parameter discovery, achieving superior QoR compared to existing tools.