### v4

# Holistic Optimization Framework for FPGA Accelerators

Abstract.
Customized accelerators have revolutionized modern computing by delivering substantial gains in energy efficiency and performance through hardware specialization. Field-Programmable Gate Arrays (FPGAs) play a crucial role in this paradigm, offering unparalleled flexibility and high-performance potential. High-Level Synthesis (HLS) and source-to-source compilers have simplified FPGA development by translating high-level programming languages into hardware descriptions enriched with directives. However, achieving high Quality of Results (QoR) remains a significant challenge, requiring intricate code transformations, strategic directive placement, and optimized data communication.
This paper presents Prometheus, a holistic optimization framework that integrates key optimizations—including task fusion, tiling, loop permutation, computation-communication overlap, and concurrent task execution—into a unified design space. By leveraging Non-Linear Programming (NLP) methodologies, Prometheus explores the optimization space under strict resource constraints, enabling automatic bitstream generation. Unlike existing frameworks, Prometheus considers interdependent transformations and dynamically balances computation and memory access.
We evaluate Prometheus across multiple benchmarks, demonstrating its ability to maximize parallelism, minimize execution stalls, and optimize data movement. The results showcase its superior performance compared to state-of-the-art FPGA optimization frameworks, highlighting its effectiveness in delivering high QoR while reducing manual tuning efforts.
High-Level Synthesis, NonLinear Programming, Compiler
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Hardware High-level and register-transfer level synthesisccs: Software and its engineering Compilers
1. Introduction
The rise of customized accelerators has revolutionized modern computing, driving significant improvements in energy efficiency and performance through hardware specialization. Among these accelerators, Field-Programmable Gate Arrays (FPGAs) have gained widespread adoption due to their flexibility, high performance, and adaptability across diverse domains, including machine learning, scientific computing, financial modeling, and embedded systems. Unlike fixed-function hardware such as Application-Specific Integrated Circuits (ASICs), FPGAs offer reconfigurability, enabling hardware adaptation to workload-specific requirements. However, achieving a good Quality of Result (QoR) on FPGAs remains a complex and resource-intensive process, requiring careful tuning of computation, memory access, and parallelism strategies.
To address this complexity, High-Level Synthesis (HLS) tools and source-to-source compilers (Cong et al., 2011; Zhang et al., 2008; Xilinx, 0232; Intel, 2024; Siemens, 2023; Microchip, 2023; Xilinx, 2024; Huang et al., 2021; Ye et al., 2022; Lai et al., 2019; Yu et al., 2018; Li et al., 2020; Pouget et al., 2025b; Basalama and Cong, 2025) have emerged as critical enablers of FPGA adoption, allowing developers to write hardware-accelerated programs in high-level languages such as C++ and Python. These tools generate synthesizable hardware descriptions, leveraging pragmas (hardware directives) and code transformations to optimize execution. Despite their advantages, HLS-generated designs still require substantial manual tuning to achieve high QoR, as the compiler’s ability to optimize performance remains highly dependent on the structure and directives applied to the input code.
1.1. Challenges in FPGA Optimization
Despite significant advancements in HLS automation and FPGA design methodologies, achieving high-performance and resource-efficient FPGA implementations remains a challenging problem. This complexity arises from the intricate interactions between computation, memory access, and parallel execution strategies. Several key challenges must be addressed to unlock the full potential of FPGA acceleration.
Code Transformation and Design Space Exploration: FPGA performance is highly dependent on how computations are structured and scheduled. Optimizations such as loop tiling, loop permutation, and loop unrolling play a crucial role in exposing parallelism and improving memory access efficiency.
Task Concurrency and Resource Management: Maximizing throughput in FPGA designs requires efficient task concurrency, where multiple computation kernels execute in parallel. However, coordinating parallel tasks introduces challenges in managing resource contention, memory bandwidth, and routing complexity. Overlapping computations effectively demands intelligent scheduling strategies that minimize stall cycles while avoiding bottlenecks caused by limited on-chip storage and off-chip memory latency.
Computation-Communication Overlap: FPGA acceleration is often hindered by inefficient data movement between off-chip memory (DDR/HBM) and on-chip compute units. Many workloads experience severe performance degradation due to memory access bottlenecks rather than computational limitations. Dataflow architectures aim to address this by implementing FIFO-based streaming to pipeline data transfers, but they often fail to exploit intra-task parallelism efficiently. In contrast, shared buffer models (Pouget et al., 2025b; Ye et al., 2022) improve data reuse, but their rigid memory management strategies make it difficult to overlap computation and communication dynamically. A more adaptive approach is needed to orchestrate data movement in parallel with execution, reducing overall memory stalls.
Scalability and Multi-SLR Utilization: FPGA architectures, especially those using Stacked Silicon Interconnect (SSI) technology, feature multiple Super Logic Regions (SLRs) to enhance scalability. However, most frameworks, including Merlin-based tools (e.g., AutoDSE, HARP, NLP-DSE) and Sisyphus, restrict designs to a single SLR, leading to underutilization of FPGA resources. While designing across multiple SLRs is possible, it significantly increases routing congestion, timing closure failures, and bitstream generation complexity. Without an SLR-aware optimization strategy, large-scale tasks face a high probability of bitstream generation failures or degraded performance due to excessive inter-SLR communication overhead.
Comprehensive Design Space Exploration: Current design space exploration (DSE) methodologies are often narrow in scope, considering only a limited subset of optimizations at a time. Each transformation—whether it involves loop reordering, tiling, pipelining, or data partitioning—has a profound impact on the overall performance and resource usage. Frameworks that optimize only one aspect of the design fail to account for the interdependencies between different optimizations, leading to suboptimal results. To achieve maximum efficiency, FPGA designs require a global optimization approach that integrates all possible transformations while considering the constraints imposed by memory hierarchy, computation parallelism, and hardware utilization.
Addressing these challenges requires a unified optimization strategy that seamlessly integrates code transformation, task concurrency management, computation-communication overlap, and hardware-aware scheduling into a single framework. The proposed work aims to develop such an approach, ensuring high-performance FPGA implementations while maintaining synthesis feasibility.
1.2. Contributions and Prometheus Method
This work introduces Prometheus, a unified framework that automates FPGA design space exploration by integrating multiple optimizations into a single methodology. Unlike existing tools that optimize isolated aspects of FPGA acceleration, Prometheus holistically addresses code transformation, memory management, task concurrency, and hardware-aware scheduling. It leverages a hybrid execution model that balances shared memory reuse and dataflow streaming to maximize parallelism while minimizing memory overhead.
To explore the vast design space efficiently, Prometheus formulates the optimization process as a Non-Linear Programming (NLP) problem, enabling automatic selection of loop transformations, tiling strategies, pragma configurations, and memory partitioning policies. Unlike heuristic-based approaches, this formulation captures complex optimization interdependencies and ensures globally efficient designs. Additionally, Prometheus incorporates SLR-aware task scheduling, which partitions tasks across multiple Super Logic Regions (SLRs). This addresses a major limitation in prior works that restrict execution to a single SLR, leading to underutilization of FPGA resources. By integrating SLR-aware optimizations, Prometheus reduces routing congestion, improves scalability, and ensures successful bitstream generation.
The methodology follows a structured process: First, Prometheus performs affine code analysis to identify parallelism opportunities and dependencies. It then generates a dataflow graph and fuses tasks that share the same outputs to minimize unnecessary communication overhead. Next, Prometheus constructs a comprehensive design space that includes all key optimizations outlined in Section 2, such as task concurrency via dataflow, computation-communication overlap, and adaptive parallelism, while ensuring resource constraints are met for each SLR. The NLP-based design space exploration model, as described in Section 4, is then employed to identify the theoretically optimal set of transformations. Once the best configuration is identified, Prometheus automatically generates the optimized design. Finally, the system produces an OpenCL host code and compiles the FPGA bitstream with the HLS compiler.
The key contributions of this work are:
  • Unified FPGA optimization framework that jointly optimizes loop transformations, pragma insertion, task concurency and computation-communication overlap while considering interdependencies between computation and data movement.
  • Hybrid execution model that dynamically selects between shared buffering and dataflow streaming to maximize parallelism and efficient memory access.
  • NLP-based design space exploration that automates the selection of loop tiling, scheduling, and memory strategies for globally theoretical optimal performance.
  • SLR-aware scheduling and multi-SLR partitioning to balance routing complexity and FPGA resource utilization, overcoming single-SLR limitations in prior works.
  • End-to-end compilation and automation that generates optimized HLS-C++ code, OpenCL host code, and FPGA bitstreams with minimal manual intervention.
  • Comprehensive performance evaluation, demonstrating superior QoR compared to AutoDSE (Sohrabizadeh et al., 2021), Sisyphus (Pouget et al., 2025b), ScaleHLS (Ye et al., 2022), Stream-HLS (Basalama and Cong, 2025) and Allo (Chen et al., 2024).
2. Background and Motivation
Optimizing latency and performance in FPGA designs relies on a combination of code transformations and strategic insertion of directives within High-Level Synthesis (HLS) (Cong et al., 2018). This section explores various HLS optimization techniques, their role in enhancing parallelism and resource utilization, as well as their limitations and challenges.
Table 1 presents a comparative analysis of various frameworks built on top of AMD Vitis HLS. These frameworks are categorized based on their underlying methodologies, including Model-Free, AI-Based, Cost Model Communication, Cost Model Computation, and NLP-Based approaches. The table evaluates their capabilities across key optimization techniques and objectives.
Model Free AI Model Cost Model Communication Cost Model Computation NLP-Based
AutoDSE (Sohrabizadeh et al., 2021) HARP (Sohrabizadeh et al., 2023) PolyOpt-HLS (Pouchet et al., 2013a) ScaleHLS (Ye et al., 2022) / POM (Zhang et al., 2024) HeteroCL (Lai et al., 2019) / Allo (Chen et al., 2024) StreamHLS (Basalama and Cong, 2025) NLP-DSE (Pouget et al., 2025a) Sisyphus (Pouget et al., 2025b) Prometheus
Pragma Insertion Limit.
Code Transformation (Tiling) Limit Limit
Code Transformation (Loop Permutation) Limit Limit
Code transformation + pragma insertion (unified)
Task Concurrency Limit.
Dataflow
Computation-Communication Overlap Limit.
Data Packing
Padding (Communication)
Padding (Computation)
Hardware Aware Limit.
Hardware-feasible Limit. Limit. Limit. Limit.
Management of Off-Chip and On-Chip Memory Transfers
Objective Comm + Comp Comm + Comp Comm Comp - Comp Comm + Comp Comm + Comp Comm + Comp
Enumeration (AI, heuristics, …) Manual
Table 1. Comparison of HLS-Based FPGA Optimization Frameworks. This table summarizes key features supported by various frameworks, including pragma insertion, code transformations, task concurrency, dataflow modeling, and hardware-awareness, categorized by their underlying optimization strategies.
2.1. HLS Optimization
2.1.1. Pragma Insertion
HLS optimizations rely extensively on a range of directives that control loop execution, data access patterns, and computation scheduling to maximize performance. Among these, three fundamental pragmas—unroll, pipeline, and array partitioning—are particularly critical for improving parallelism and reducing execution latency.
The unroll directive enables the expansion of loop iterations in the hardware design, allowing multiple iterations to execute in parallel. By eliminating loop control overhead and enabling greater resource utilization, unrolling can significantly boost throughput. However, its effectiveness is limited by FPGA resource constraints, such as available DSP slices, BRAM, and routing congestion.
The pipeline pragma restructures loop execution to allow overlapping operations, thereby reducing the initiation interval (II)—the number of cycles required to launch consecutive iterations. This directive is essential for achieving high throughput in applications with iterative computations, such as matrix multiplications and stencil operations. Careful tuning of pipeline depth and initiation interval is necessary to balance resource usage while avoiding performance bottlenecks due to memory access contention.
Array partitioning is another crucial optimization that enhances memory access parallelism. By splitting large arrays into smaller banks stored in separate on-chip BRAM blocks or distributed memory structures, this pragma allows simultaneous data accesses, facilitating more effective unrolling and pipelining. Without partitioning, memory conflicts and bandwidth limitations can restrict performance, especially with multiple concurrent memory accesses.
Beyond these fundamental pragmas, additional directives such as loop flattening, and dependency pragmas can further optimize execution. Loop flattening combines nested loops into a single loop to improve hardware efficiency. Dependency pragmas help manage data dependencies to ensure efficient scheduling without unnecessary synchronization delays.
Pragma insertion has been extensively studied using various methodologies to optimize performance and resource utilization in HLS designs. These approaches include bottleneck analysis, as employed by AutoDSE (Sohrabizadeh et al., 2021), performance estimation through Graph Neural Networks (GNN) in frameworks like Harp (Sohrabizadeh et al., 2023), and cost model optimization leveraging Non-Linear Programming (NLP) in NLP-DSE (Pouget et al., 2025a, 2024).
Each method presents distinct advantages and limitations. AutoDSE achieves full accuracy by running the HLS compiler for each configuration, ensuring precise performance measurements. However, this approach is computationally expensive and time-consuming, making it impractical for rapid design exploration. HARP, on the other hand, can estimate performance across configurations without direct compilation, significantly reducing evaluation time. However, it requires extensive training datasets and fine-tuning for each new kernel to accurately model compiler behavior. While this approach enables learning-based adaptation to different designs, its effectiveness depends heavily on data quality and model calibration. Conversely, NLP-DSE explores the entire design space efficiently by formulating the optimization as a Non-Linear Programming problem. This method allows for rapid exploration and selection of theoretical optimal configurations. However, it relies on a theoretical cost model, which may introduce inaccuracies if the compiler behaves unpredictably or if certain optimizations are not accurately captured by the model. Consequently, while NLP-DSE provides a fast and scalable solution, its reliability depends on the fidelity of the cost model to actual HLS compilation behavior.
While pragma insertion is a powerful optimization technique, its effectiveness is inherently dependent on the interaction between memory hierarchy, computational resources, and FPGA-specific constraints. Moreover, its impact is tightly coupled with the program’s execution schedule, meaning that without appropriate code transformations, the benefits of pragma insertion remain limited. Effective optimization requires a synergy between pragma directives and structural modifications to the code to fully exploit FPGA parallelism and resource utilization.
2.1.2. Code Transformation
Code transformations are fundamental for optimizing execution on FPGAs. While pragma insertion plays a crucial role in enhancing performance, it must be complemented by effective code transformations to fully exploit FPGA resources. Techniques such as loop reordering (permutation), task fusion, and data tiling enhance data locality and increase parallelism, thereby improving both computational efficiency and memory access patterns.
Various code transformation strategies have been developed specifically for FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). While transformations originally designed for CPUs and GPUs achieve substantial performance gains by optimizing for their respective architectures, they do not inherently align with FPGA requirements, which prioritize fine-grained parallelism and efficient resource utilization. Several studies have utilized Pluto (Bondhugula et al., 2008), a leading compiler framework originally designed for CPU optimizations, to transform FPGA kernels (Zhao et al., 2022; Zhao and Cheng, 2021). While Pluto excels at tiling and minimizing dependencies to enhance memory reuse, its direct application to FPGA optimization is constrained due to the fundamental differences in optimization strategies required for CPUs and FPGAs. Unlike CPUs, where memory hierarchy and cache locality are primary concerns, FPGA optimization focuses on maximizing parallelism, minimizing resource contention, and efficiently utilizing on-chip memory, making Pluto’s conventional transformation techniques less effective in this context. Conversely, studies such as (Pouchet et al., 2013b; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on code transformations tailored to specific FPGA performance goals. The work in (Pouchet et al., 2013b) aims to reduce communication overhead between off-chip and on-chip memory, achieving superior Quality of Results (QoR) for memory-bound kernels. Meanwhile, research efforts in (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) concentrate on optimizing pipelining strategies to maximize instruction-level parallelism and resource utilization. Sisyphus (Pouget et al., 2025b) introduces a unified approach by integrating code transformation and pragma insertion into a single optimization problem. By formulating this as a Non-Linear Programming (NLP) problem, Sisyphus efficiently explores the design space to identify theoretical optimal configurations, streamlining FPGA acceleration while maintaining a balance between computation and memory access efficiency.
2.1.3. Task Concurrency
HLS tools like Vitis HLS support the dataflow pragma, which structures computations into actors that communicate through FIFO queues. This approach allows overlapping execution of multiple tasks, significantly reducing overall latency . By enforcing a producer-consumer model, dataflow scheduling enables each task to process data as soon as it becomes available, rather than waiting for an entire stage to complete, which is particularly beneficial.
For computational kernels such as 3mm (Listing 3), the dataflow paradigm allows the first two matrix multiplications to execute in parallel while the third begins as soon as its required inputs are produced. This overlapping of execution helps maximize throughput and minimize idle time for computing units. Additionally, dataflow optimizations facilitate task-level parallelism, allowing independent tasks to run concurrently across multiple compute resources, such as DSP blocks and BRAM, ensuring efficient utilization of available FPGA resources.
However, despite its advantages, pure dataflow parallelism presents several challenges. One major limitation is intra-task parallelism, which is constrained by the reliance on FIFO-based communication. Since each FIFO can only transfer up to 512 bits per cycle (as this is the maximum off-chip memory bitwidth supported), this inherently restricts the amount of data that can be processed concurrently within a task. To further enhance intra-task parallelism, alternative approaches must be explored.
Stream-HLS (Basalama and Cong, 2025) attempts to address this limitation by increasing the number of FIFOs connecting two tasks, assuming that all data resides on-chip. While this approach simplifies certain aspects of the design space, it is neither scalable nor generalizable for transferring data from off-chip to on-chip memory. By using $n$ FIFOs to transfer $n$ data elements in parallel between two tasks, the method significantly increases resource consumption without proportionally improving efficiency. The additional FIFOs complicate the design, potentially leading to routing congestion and excessive hardware overhead.
Moreover, modern FPGAs feature up to 32 off-chip memory banks, making the multi-FIFO approach impractical for handling off-chip data transfers. A more effective and scalable strategy must be developed to optimize intra-task parallelism while ensuring efficient communication between off-chip and on-chip memory, without introducing unnecessary complexity or resource constraints.
2.1.4. Shared Buffering
Shared buffering is a critical technique in FPGA memory optimization, enabling efficient data access and reuse across multiple computational units. It can be employed at a global level, as seen in frameworks like AutoDSE (Sohrabizadeh et al., 2021), Sisyphus (Pouget et al., 2025b), and ScaleHLS (Ye et al., 2022), or within individual dataflow tasks to enhance execution efficiency.
This approach involves preloading data buffers into on-chip memory, allowing multiple computations to access shared data without redundant transfers. By reducing memory access latency and improving bandwidth utilization, shared buffering facilitates high parallelism and optimizes overall performance. It is particularly beneficial in applications which are computation bound.
However, shared buffering presents challenges in maintaining concurrency. While it enables efficient data reuse, it can introduce synchronization overhead, especially when multiple tasks attempt to access the same memory region simultaneously. Managing concurrent access requires arbitration mechanisms, which can lead to increased latency and potential bottlenecks. Additionally, routing congestion can occur due to high interconnect demands, limiting scalability in complex FPGA designs.
One limitation of shared buffering is its impact on initiation interval (II) in pipelined architectures. Unlike dataflow-based designs that rely on FIFO queues for seamless data streaming, shared buffering requires explicit read and write coordination, which may introduce stalls if not carefully managed. Moreover, FPGA resource constraints, such as limited BRAM and URAM availability, impose restrictions on buffer size and allocation strategies.
To address these challenges, advanced techniques such as double-buffering, memory partitioning, and adaptive scheduling have been explored. Double-buffering enables overlapping computation with data transfer, reducing idle cycles. Memory partitioning distributes data across multiple banks to alleviate contention, while adaptive scheduling dynamically assigns buffer access based on task priority and workload demands.
By effectively integrating shared buffering with intelligent memory management strategies, FPGA designs can achieve a balance between computational parallelism and efficient memory access. Future advancements should focus on automated buffer allocation techniques and dynamic access pattern optimization to further enhance performance in high-level synthesis (HLS) workflows.
2.1.5. Computation-Communication Overlap
Overlapping computation and communication is crucial for high-performance FPGA designs. Techniques such as double buffering (ping-pong buffering) and advanced data tiling help mask communication latency while keeping computational units busy. Managing data transfers efficiently between on-chip and off-chip memory ensures that processing units remain active without waiting for data.
2.1.6. Data Packing and Padding
Modern FPGA architectures support high-bandwidth data transfers (up to 512-bit wide on AMD/Xilinx FPGAs). Data packing and padding optimize memory alignment, reducing the number of required memory cycles. For instance, transferring 216 floating-point values using a 256-bit width (8 floats per transfer) requires 27 cycles, making efficient packing crucial for minimizing overhead.
To fully exploit data packing, padding must be considered to enable even faster data transfers. Additionally, padding is valuable for achieving finer control over parallelism and resource utilization by expanding the available design space for loop unrolling, thereby enhancing computational throughput and efficiency.
Padding for Communication
Padding must be considered to increase the flexibility of data packing. The original size of the data may impose restrictions on packing efficiency, but by introducing padding, we create a larger space that may allow for a more efficient transfer. However, padding is not a free optimization, as increasing it also increases the amount of data that needs to be transferred.
For example, in the code presented in Listings 1 and 2, where $J=190$, transferring all elements along the second dimension (iterated by loop $j$) of array $B$ onto the chip is constrained by the available memory bandwidth. Without padding, the maximum transfer rate is $64$ bits per cycle, as $190\times 32$ is divisible by $64$ but not by $128$. However, by introducing padding with $P=2$ and adjusting $J$ to $192$ ($190+2$), the data alignment enables a significantly higher transfer rate of $512$ bits per cycle, as $192\times 32$ is now perfectly divisible by $512$. This optimization maximizes memory bandwidth utilization, reducing transfer latency and improving overall data throughput.
Padding for Computation
Padding can also be used for computation, if we take a similar method than Sisyphus (Pouget et al., 2025b) which tiled and unroll the intra-tile loops which correspond to the transformation from Listing 1 to Listing 2. We will have an unroll factor which correspond to $I1\times J1\times K1$ but we want to have $I1$ divide $I$, $J1$ divide $J$ and $K1$ divide K because we do not want to use extra ressource to only compute a partial tile which would correspond to the execution of the last tile which is not complete. For example if we have a trip count of $190$ and we use an unroll factor of $8$ then we will execute $184\times 8$ a full tile and then 6 iterations for the partial tile.
But we restrict the unroll factor to be divisible of the trip count we restrict considerably the space of possibilities. So the solution is to pad which allow to considerably increase the space of possibilities. Hence with a trip count of $190$ the possible unroll factor are include in $UF\in\{1,2,5,10,19,38,95,190\}$ but with padding of 2 (a trip count of 192) the space become $UF\in\{1,2,3,4,6,8,12,16,24,32,48,64,96,192\}$. And of course all pading can be considered. Padding enables finer control over unrolling and resource utilization, allowing for more flexible and efficient hardware implementation while optimizing computational parallelism.
1for (i = 0; i < I; i++)
2 for (j = 0; j < J; j++)
3 for (k = 0; k < K; k++)
4 C[i][j] += A[i][k] * B[k][j];
Listing 1: Baseline Implementation of Matrix Multiplication in C. This code depicts a naive triple-nested loop structure used for matrix-matrix multiplication, serving as the unoptimized reference for further transformations.
1for (i0 = 0; i0 < I0; i0++)
2 for (j0 = 0; j0 < J0; j0++)
3 for (k0 = 0; k0 < K0; k0++)
4 for (i1 = 0; i1 < I1; i1++)
5#pragma HLS unroll
6 for (j1 = 0; j1 < J1; j1++)
7#pragma HLS unroll
8 for (k1 = 0; k1 < K1; k1++)
9#pragma HLS unroll
10 C[i0*I1 + i1][j0*J1 + j1] += A[i0*I1 + i1][k0*K1 + k1] * B[k0*K1 + k1][j0*J1 + j1];
Listing 2: Matrix Multiplication with Loop Tiling and Fully Unrolled Intra-Tile Computation. This implementation showcases a performance-optimized matrix multiplication using loop tiling to enhance data locality and fully unrolled intra-tile loops to expose fine-grained parallelism.
2.2. Performance and Hardware Considerations
2.2.1. Performance Evaluation
Performance evaluation in HLS is a multi-stage process that assesses different aspects of a design’s efficiency and feasibility. It is typically conducted at three levels: estimation (e.g., Vitis HLS reports), RTL simulation, and FPGA-based evaluation. Each of these stages provides critical insights, but they vary in accuracy, speed, and resource requirements.
  1. (1)
    Estimation (HLS Reports): The first level of evaluation relies on estimation tools provided by HLS compilers such as Vitis HLS. These reports generate quick approximations of key performance metrics, including latency, resource utilization (DSPs, LUTs, BRAMs), and throughput. While these estimations are useful for early-stage design exploration, they do not account for low-level placement and routing effects, which can significantly impact real-world performance. The optimistic assumptions of HLS reports often fail to capture routing congestion, memory access delays, and other architectural bottlenecks.
  2. (2)
    RTL Simulation: To achieve more accurate performance insights, RTL simulation is performed after HLS compilation. This stage involves generating synthesizable RTL code, which is then simulated using hardware description language (HDL) tools. RTL simulation provides cycle-accurate performance metrics, allowing designers to analyze pipeline behavior, data dependencies, and execution timing. However, while RTL simulation models hardware more precisely than HLS reports, it does not incorporate real-world FPGA constraints such as clock tree synthesis, interconnect delays, and power distribution, which can affect the final implementation.
  3. (3)
    FPGA-Based Evaluation: The most accurate method of performance evaluation involves deploying the design onto an actual FPGA. This process includes place-and-route, bitstream generation, and execution on hardware. FPGA-based evaluation provides real execution metrics such as operating frequency, power consumption, and effective memory bandwidth. It also reveals practical challenges, including timing closure issues, placement inefficiencies, and resource contention that cannot be detected in earlier stages. This level of evaluation is essential for validating Quality of Results (QoR) and ensuring the design meets real-world constraints.
The transition from RTL simulation to FPGA implementation often introduces additional challenges, such as increased routing complexity, unexpected timing violations, and suboptimal resource utilization. These discrepancies arise due to the abstraction gap between HLS-generated code and final FPGA hardware constraints. To mitigate these issues, it is essential to develop robust tools that enable rapid design regeneration by modifying specific configurations within a well-defined portion of the design space. Such tools would enable engineers to efficiently iterate on design parameters without requiring a complete code regeneration, which can be time-consuming and may result in a significantly different design that still fails to generate a valid bitstream. Having the ability to selectively modify only the congested parts of the design while preserving the rest of the configuration would be highly valuable, ensuring faster convergence toward an optimized and feasible FPGA implementation.
2.2.2. Hardware Awareness and Resource Constraints
Many existing studies conclude their evaluations at the Vitis HLS report or RTL simulation stage, overlooking the critical impact of placement and routing constraints. However, as the design progresses toward bitstream generation, the available design space becomes increasingly constrained. Addressing hardware limitations early in the HLS process is essential to avoid costly design iterations and ensure convergence toward a feasible and deployable FPGA implementation.
Additionally, Super Logic Region (SLR)-aware optimizations are crucial for multi-SLR FPGA architectures. A Super Logic Region (SLR) is a physically distinct section of an FPGA die in Stacked Silicon Interconnect (SSI) technology. Each SLR contains a subset of the device’s computational resources, including DSP, LUT and FF, similar to those in monolithic FPGA devices.
Effectively mapping tasks across SLRs is critical for balancing resource utilization and minimizing routing congestion, as inter-SLR communication introduces additional latency and can become a performance bottleneck. By integrating SLR-aware task partitioning, optimized scheduling, and efficient data movement strategies, FPGA designs can achieve higher scalability, improved parallelism, and enhanced synthesis feasibility. Ensuring that computation and memory access patterns align with the physical structure of multi-SLR FPGAs is essential for achieving high QoR while maintaining timing closure and efficient resource allocation.
Some recent frameworks, such as RapidStream-TAPA (Guo et al., 2023a, b) and PASTA (Khatti et al., 2023), aim to improve the design scalability and performance of HLS programs on modern multi-die FPGAs through hardware-aware co-optimization of HLS and physical design. While RapidStream-TAPA focuses on latency-insensitive, FIFO-based communication and exploits coarse-grained floorplanning and pipelining for timing closure and parallel compilation, PASTA extends this model to include buffer-based inter-task communication using a generalized channel abstraction. Despite their advances, both frameworks still require a manually optimized task-parallel input code, which demands substantial expertise in HLS and hardware design, limiting their accessibility for non-expert users.
2.3. Challenges
While existing HLS optimizations provide substantial performance improvements, several challenges remain. One of the primary limitations is that many prior works restrict their optimization space to specific techniques or separate their optimization processes into multiple independent steps. This fragmented approach can lead to incoherent design decisions, where optimizations applied at one stage may not align with or may even counteract those applied at subsequent stages. Refer to Table 1 for a detailed comparison of these approaches.
A common oversimplification in many frameworks is the assumption that all data resides on-chip. While this simplifies memory access patterns in design space exploration, it often results in low Quality of Results (QoR) when deployed on real hardware. This assumption prevents frameworks from incorporating crucial tiling strategies, which are essential for optimizing data locality, reducing memory access overhead, and ensuring efficient off-chip communication. When off-chip memory management is treated as a separate optimization step, decisions made in earlier phases may restrict the effectiveness of later memory optimizations, leading to suboptimal performance and resource utilization.
The limitations of current HLS optimization frameworks that rely on shared buffering are exemplified through the 3mm kernel. It consists of two independent matrix multiplications, whose outputs serve as inputs for a final matrix multiplication.
As illustrated in Listing 3, each matrix multiplication consists of deeply nested loops that are well-suited to parallelization techniques such as loop unrolling, pipelining, and computation-communication overlap. However, conventional shared buffering strategies often fall short in effectively leveraging concurrency, primarily due to their limited exploitation of dataflow principles and insufficient overlap between computation and communication.
Another major challenge is balancing parallelism and routing complexity. Excessive parallelization can lead to routing congestion, increased place-and-route time, and ultimately, failure to generate a valid bitstream. Current methodologies lack adaptive mechanisms to dynamically adjust parallelism levels based on available routing resources and FPGA architecture constraints. Without such mechanisms, achieving high-performance designs often requires manual intervention and iterative fine-tuning.
While pragma-based optimizations such as unroll and pipeline significantly improve performance, their effectiveness is inherently tied to the underlying code structure. Without proper transformations to expose parallelism and optimize memory access patterns, pragma insertion alone cannot fully unlock the potential of FPGA acceleration.
Addressing these challenges requires a holistic approach that integrates memory-aware optimizations, automated design space exploration, and adaptive scheduling techniques. Future research should focus on frameworks that jointly optimize data placement, computational parallelism, and resource utilization while considering real hardware constraints. By developing more cohesive and intelligent HLS optimization strategies, FPGA-based acceleration can achieve greater efficiency, scalability, and deployment feasibility.
Model-free approaches, such as AutoDSE (Sohrabizadeh et al., 2021) and AI-based frameworks like HARP (Sohrabizadeh et al., 2023; Wu et al., 2023, 2022a), rely on either HLS compilers or AI models to predict performance and resource utilization. However, these methods are currently limited to pragma insertion and space enumeration. Even though AI-based models provide fast estimations (on the order of milliseconds), exhaustively exploring large design spaces remains infeasible. Our previous work, NLP-DSE (Pouget et al., 2025a), addresses this limitation by leveraging NLP techniques to efficiently explore pragma configurations. However, like the other approaches, it remains restricted to pragma insertion.
PolyOptHLS (Pouchet et al., 2013a) focuses on optimizing designs by minimizing memory transfers. While reducing communication latency can significantly improve performance, it does not always lead to optimal results. In many cases, the primary bottleneck may shift from communication to computation, limiting overall efficiency and potentially degrading the quality of results (QoR). This approach, therefore, lacks a balanced optimization strategy that considers both computation and communication trade-offs.
ScaleHLS (Ye et al., 2022) and POM (Zhang et al., 2024) extend optimization beyond pragma insertion by incorporating code transformations. However, their exploration space is constrained by the assumption that data is already on-chip. Additionally, their transformations are based on heuristics, such as permuting the reduction loop to the outermost level. While they employ a cost model to minimize computation latency, they still rely on exhaustive enumeration of possible configurations.
Allo (Chen et al., 2024) aims to reduce development time and enhance design quality. However, it still requires significant manual intervention from an expert, limiting its usability for non-specialists.
Our previous work, Sisyphus (Pouget et al., 2025b), enables both code transformations and pragma insertion but is restricted to optimizing a single task. It lacks support for optimizations that dataflow techniques can provide, which would be necessary for broader efficiency improvements.
Stream-HLS (Basalama and Cong, 2025) effectively leverages dataflow pragmas by selecting a good loop ordering strategy to maximize streaming efficiency. However, it imposes constraints on the design space by assuming that data is on-chip. Moreover, its approach to increasing parallelism relies on multiple FIFOs, which is not generalizable when off-chip memory access is required.
A significant limitation of all these frameworks is that none of them account for hardware constraints in their exploration space. While this simplifies the exploration process, it reduces real-world applicability. Additionally, frameworks such as Merlin-based tools (AutoDSE, HARP, NLP-DSE) and Sisyphus can generate all necessary components, including off-chip memory, for bitstream generation. However, they are constrained to a single Super Logic Region (SLR), leading to under-utilization of the FPGA board. Although generating a design that spans multiple SLRs is theoretically possible, it significantly increases the risk of bitstream generation failures. Even if the bitstream is successfully generated, timing violations often degrade performance.
1for (i = 0; i < 180; i++) // MM 1
2 for (j = 0; j < 190; j++) {
3 E[i][j] = 0.0; // S0
4 for (k = 0; k < 200; ++k)
5 E[i][j] += A[i][k] * B[k][j];} // S1
6for (i = 0; i < 190; i++) // MM 2
7 for (j = 0; j < 210; j++) {
8 F[i][j] = 0.0; // S2
9 for (k = 0; k < 220; ++k)
10 F[i][j] += C[i][k] * D[k][j];} // S3
11for (i = 0; i < 180; i++) // MM 3
12 for (j = 0; j < 210; j++) {
13 G[i][j] = 0.0; // S4
14 for (k = 0; k < 190; ++k)
15 G[i][j] += E[i][k] * F[k][j];} // S5
Listing 3: Reference Implementation of the 3mm Kernel from PolyBench. This listing shows the baseline version of the 3mm benchmark, comprising three matrix multiplication operations. The first two multiplications are independent and produce intermediate results that serve as inputs to the final matrix multiplication.
2.4. Overview of Prometheus
We introduce Prometheus, a unified optimization framework designed to efficiently navigate complex design spaces using a Non-Linear Programming (NLP) cost model. Prometheus streamlines FPGA design development by providing an intuitive, engineer-friendly interface with tunable parameters, including tile size, array partitioning, and loop unrolling factors.
The framework integrates all the optimization techniques discussed in previous sections, ensuring a comprehensive and cohesive approach to performance enhancement. Additionally, Prometheus is SLR-aware, enabling intelligent task distribution across Super Logic Regions (SLRs). This capability not only simplifies bitstream generation but also optimizes the trade-off between computational performance and resource utilization.
Prometheus C++ Code Intermediate Representation (dependencies, trip counts) Task-flow Graph
1234
Formulate a Non-Linear Problem encoding the full optimization design space All Parameters of the Spacee.g. data-tile of array $A$ is $\texttt{32}\times\texttt{64}$Files for HLS CompilerHLS-C++, OpenCL, Config Bitstream with HLS Compiler
Extract with PoCC GenerateSolve NLP Generate
Figure 1. Overview of the Prometheus Framework Workflow. This diagram illustrates the end-to-end pipeline of the Prometheus optimization framework, starting from C++ source code and proceeding through intermediate representation extraction, task-flow graph generation, Non-Linear Programming (NLP)-based design space exploration, and final compilation into FPGA bitstreams using HLS compilers.
Group Parameter Description
Design Variables
Memory Bit Width ($BW_{a}$) Width of memory transfer in bits;
Data-tile Size (modeled via Transfer Level $t_{a,l}$ in Section 4) Size of the data tile of array $a$ transferred either from off-chip to on-chip memory or between tasks. The transfer level determines this size.
Data-tile Reuse Size (modeled via Reuse Level $d_{a,l}$ in Section 4) Size of the data tile of array $a$ reused within a task. The reuse level dictates this tile size.
Buffering ($N_{a}$) Number of buffers (2 for double buffering, 3 if read & write), enabling overlap of load/compute/store.
Communication Padding Additional data added to enable wider burst transfers and better bandwidth utilization.
Parallelism Unroll Factors ($TC^{l}_{\text{intra}}$) Unroll factor for each task, determined through tiling, where the intra-tile loop is fully unrolled to expose fine-grained parallelism.
Array Partitioning ($AP_{a,d}$) Number of partitions for array $a$ along dimension $d$ to support parallel accesses.
Compute Padding Padding applied to loop trip counts to increase legal unroll factors.
Code Structure Loop Permutations Legal reordering of loops to optimize data locality and parallelism; synchronized across fused statements.
Tiling Splitting loops into inter- and intra-tile levels for optimization and pipelining.
SLR Assignment ($slr_{t}$) Mapping of each task $t$ to a Super Logic Region (SLR) to enable spatial task distribution.
Design Constraints
Resource Limits Max Array Partitioning Total partitions across all arrays must not exceed architectural limits.
DSP Budget Maximum DSP slices per SLR; used to limit total DSP usage by pipelined/unrolled tasks.
On-Chip Memory Total memory usage by buffered tiles (considering reuse and buffering) must be within BRAM capacity.
Table 2. Design Variables and Architectural Constraints in the Prometheus Optimization Space. This table outlines the key configurable parameters used in Prometheus—including loop tiling, unrolling, array partitioning, and buffer management—along with hardware constraints such as resource limits and memory capacity. The cost model also accounts for concurrent task execution enabled by the dataflow pragma, as well as computation-communication overlap achieved through automatic double or triple buffering. These elements jointly define the valid and efficient design space explored by Prometheus.
Figure 1 illustrates the workflow of Prometheus. The process begins with an input C++ code, from which we extract an intermediate representation containing dependencies, trip counts, schedules, and other relevant information. This extraction is performed using PoCC (Pouchet, 2023), which provides all necessary details for analysis. Next, we construct the task-flow graph based on this extracted information and formulate the corresponding Non-Linear Programming (NLP) problem. Solving this NLP problem determines the theoretical optimal parameters for the design space. Once the NLP solver provides these parameters, we have all the required information to proceed. At this stage, we automatically generate the HLS-C++ code along with all the necessary files to produce the FPGA bitstream, ensuring an efficient and fully automated compilation process.
To support efficient exploration of the design space, Prometheus defines a set of tunable parameters and architectural constraints that capture the key aspects of FPGA acceleration. Table 2 provides a detailed overview of these parameters, including memory-related configurations (e.g., bitwidth, tiling, and reuse levels), parallelism controls (e.g., loop unrolling and array partitioning), and structural transformations (e.g., loop permutations and SLR assignments). In addition, the table outlines hardware constraints such as DSP budgets, on-chip memory limits, and maximum legal array partitioning. These variables and constraints jointly define the feasible design space that the NLP-based optimization engine systematically explores to generate high-performance, hardware-aware FPGA implementations.
The pseudo-code in Listing 4 demonstrates how Prometheus processes the 3mm kernel (Listing 3). Load and store operations manage data transfers to and from off-chip memory, while send and receive operations handle inter-task communication using FIFO.
Each taski in the pseudo-code corresponds to the fully unrolled computation of an intra-tile for statement Si in Listing 3. An example of such a task is illustrated in Listing 5.
Arrays $E$, $F$, and $G$ are initialized to zero within their respective tasks (e.g., S0, S2, and S4) and are not preloaded, reducing unnecessary memory overhead.
Prometheus enhances computation efficiency by fusing statements that produce the same outputs (e.g., in Listing 3), and it automatically formulates an NLP problem to determine the theoretically optimal parameters, such as loop schedules, array bit widths (e.g., 512 bits), tile sizes, reuse buffer sizes, transfer locations, and padding. The framework dynamically adjusts bit widths for arrays like $F$, $D$, and $G$ to achieve a better balance between parallelism and resource usage.
1/*************** Fused Task 0 ****************/
2float B[204][192]; load_B(B);
3for (i0 = 0; i0 < 18; i0++)//inter-tile loop
4 float A[10][204]; load_A(A);
5 for (j0 = 0; j0 < 6; j0++){//inter-tile loop
6 float E[10][32];
7 task0(E); // S0 + intra-tile loops
8 for (k0 = 0; k0 < 51; ++k0)//inter-tile loop
9#pragma HLS pipeline II=3
10 task1(E, A, B); // S1 + intra-tile loops
11 store_E(E); sent_E(E);
12/*************** Fused Task 1 ****************/
13float C[190][222]; load_C(C);
14for (j0 = 0; j0 < 7; j0++)//inter-tile loop
15 float D[222][32]; load_D(D);
16 for (i0 = 0; i0 < 10; i0++){//inter-tile loop
17 float F[19][32];
18 task2(F); // S2 + intra-tile loops
19 for (k0 = 0; k0 < 74; ++k0)//inter-tile loop
20#pragma HLS pipeline II=3
21 task3(F, C, D); // S3 + intra-tile loops
22 store_F(F); sent_F(F);
23/*************** Fused Task 2 ****************/
24float E[180][192];
25for (j0 = 0; j0 < 7; j0++)//inter-tile loop
26 float F[192][32];
27 receive_F(F);
28 for (i0 = 0; i0 < 18; i0++){//inter-tile loop
29 float G[10][32];
30 receive_E(E);
31 task4(G); // S4 + intra-tile loops
32 for (k0 = 0; k0 < 32; ++k0)//inter-tile loop
33#pragma HLS pipeline II=3
34 task5(G, E, F); // S5 + intra-tile loops
35 store_G(G);
Listing 4: Transformed and Fused 3mm Kernel Pseudocode Generated by Prometheus. This code highlights the result of Prometheus optimizations, including fused task generation, tiled loops, memory buffer management, and pipelined execution with dataflow pragmas.
Efficient computation-communication overlap is achieved through ping-pong buffering, while concurrent task execution and FIFO-triggered dependent tasks maximize overall performance. Fused Tasks 0 and 1 execute concurrently, while Fused Task 2 begins as soon as the data tiles for $F$ and $E$ become available.
To assess the effectiveness of our proposed framework, we evaluate the performance of the 3mm kernel using several state-of-the-art HLS-based FPGA optimization tools. As shown in Table 3, Prometheus significantly outperforms existing frameworks in terms of throughput, achieving 368.36 GF/s. This represents a substantial improvement over Sisyphus (178.97 GF/s) and Stream-HLS (174.00 GF/s), and far exceeds the results of Allo, ScaleHLS, and AutoDSE. The superior performance of Prometheus stems from its holistic design space exploration strategy, which integrates loop transformations, memory tiling, concurrent task execution, and SLR-aware scheduling to deliver optimized and hardware-feasible designs.
Metric Prometheus Sisyphus Stream-HLS Allo ScaleHLS AutoDSE
Throughput (GF/s) 368.36 178.97 174.00 60.40 43.04 1.74
Table 3. Measured Throughput of the 3mm Kernel Using Various FPGA Frameworks. This table presents the runtime throughput (in GF/s) obtained from RTL simulation of the 3mm kernel across multiple frameworks. Prometheus is shown to outperform other methods, including Sisyphus, Stream-HLS, and AutoDSE.
3. Code Transformation
In our dataflow model, we adopt a synchronous dataflow where the sizes of the arrays are known during compile time. This compile-time awareness enables us to construct a precise model that facilitates rigorous optimizations. To leverage FPGA parallelism, we implement an acyclic dataflow graph, ensuring parent nodes do not receive data from their children. While this constraint limits graph configurations and may increase resource usage, it reduces overall latency. To support this structure, we inline (Chen et al., 1993) each function in the input code to generate the required acyclic graph. Our primary objective is to minimize latency within resource constraints by overlapping communication and computation within tasks and executing independent tasks concurrently. To achieve this, we apply various transformations and optimizations, explored in this section, and navigate the design space using an NLP-based approach, detailed in Section 4.
3.1. Dependency Graph Creation
Our process starts with affine C/C++ code as input, which undergoes maximal distribution to ensure each loop body contains only one statement, provided no dependencies exist within the loop. ISCC (Verdoolaege, 2011) verifies the legality of these transformations, ensuring dependencies are preserved. After achieving full distribution, we construct a dependency graph using PoCC (PoCC, [n. d.]). PoCC provides the necessary information about the schedule and dependencies, enabling us to build the graph. In this graph, the nodes represent tasks, while the edges capture data communication arising from inter-task dependencies. Tasks with identical outputs are then merged (when legal), creating fused tasks with output-stationary properties. This ensures that each tile’s output is handled (loaded, computed, and either stored or transmitted) only once. If a dependency prevents distribution, the framework will still function but with a more limited optimization space, making it more effective when distribution is possible.
Prometheus’s solution space includes:
3.2. Data-tiling and Padding
In the fused dependency graph, edges represent data communication between tasks as well as between tasks and off-chip memory, involving the transfer of data tiles with specified sizes. Data tiling divides loop iterations into smaller tiles, splitting array accesses into subsets. Each loop $l$ iterating over an array $a$ is divided into an outer loop ($TC^{l}_{inter}$) and an inner loop ($TC^{l}_{intra}$), with feasible permutations applied. For each array and dimension iterated by the loop $l$, the tile factor is a common choice that influences all arrays iterated by this loop.
To optimize memory transfers, padding is applied to arrays in two ways. Simple padding increases the bit width ($BW_{a}$) for efficient transfers while maintaining the original loop trip count. Composite padding adjusts both $BW_{a}$ and the loop trip count ($TC^{l}$) to support unroll factors that do not evenly divide the original trip count, allowing irregular tile sizes and expanding the design space. Tile sizes are consistent within a fused task but vary between tasks. In Listing 4, the tile size of array F is $19\times 32$ in Fused Task 1 and $192\times 32$ in Fused Task 2.
3.3. Fine-grained Parallelism
Data-tile selection involves splitting each loop and permuting them to create two levels of the original loops. Using ISCC (Verdoolaege, 2011), we verify the legality of these permutations, resulting in two loop levels: inter-tile (outer) and intra-tile (inner). If permutation is not feasible, the inter-tile loop retains its legal position. For tasks belonging to the same fused task, we merge their inter-tile loops, which are non-reduction. For instance, in Listing 4, the inter-tile loops on lines 3 and 5 iterate over task0 and task1. The intra-tile loops are fully unrolled, ensuring that all data accesses within the intra-tile remain on-chip. Array partitioning, determined by the unroll factor, ensures data resides in separate BRAM banks for simultaneous access. Reduction loops across inter-tiles are pipelined with an initiation interval ($II=n>1$), where n matches the reduction latency. For instance, in Listing 4, additions take 3 cycles, resulting in $II=3$.
3.4. Loop Order
Due to the full unrolling of the intra-task, there is no need to select the loop order within the intra-tile. We place the inter-tile reduction loops directly above the task and are pipelined. If multiple reduction loops exist, we rank them by the size of their trip counts, placing the loop with the highest trip count innermost to ensure an efficient pipeline. This setup provides the flexibility to choose the order of the inter-tile loops that are not reduction loops. This order will subsequently be determined by the NLP (cf. Section 4). As shown in Listing 4, the loop order of Fused Tasks 1 and 2 is permuted compared to the original program in Listing 3.
3.5. Automatic Overlapping of Communication and Computation
To overlap communication and computation, we use on-chip buffers. The buffer size and data transfer location are determined by various options explored via NLP (cf. Section 4). Since data must stay on-chip for intra-task computation, transfers occur either below an inter-tile loop or before any loops start. The transfer position determines the data tile size, covering all data accessed below it. To improve data reuse, buffer size can match or exceed the transferred data tile. Similar to transfer location, buffer size is determined by its position relative to inter-tile loops or before any loops. Two boolean variables, $d_{a,l}$ and $t_{a,l}$, define and transfer array $a$ under loop $l$ when set to true. If defined or transferred before loops, $l=0$. Double-buffering is used for read-only or write-only arrays, and triple-buffering for arrays that are both read and written. Following (Cong et al., 2018), we perform an initial load, then overlap loading the next tile with computing the current one.
3.6. Coarse-Grained Parallelism: Automatic Execution of Concurrent Tasks
Due to the use of the dataflow pragma, tasks can begin computation as soon as they have sufficient data, allowing for concurrent execution. This concurrency significantly increases overall performance.
3.7. Memory Transfer
Communication latency is reduced by transferring more data per cycle using increased bit width with padding. Read-only arrays are duplicated in off-chip memory for tasks with multiple reads, eliminating feed-through logic. For non-read-only arrays, data passes between tasks via FIFOs, using the same bit width as for off-chip memory transfer.
4. NLP Formulation
To determine the tile size, loop order, bit width, and memory transfers, we formulate a cost model as a NLP problem aimed at minimizing overall latency. This approach builds upon the methodology proposed in our previous work (Pouget et al., 2025a, 2024, b), which we have further adapted to meet the specific requirements and constraints of our current framework. In our work, we incorporate dataflow considerations along with all the optimizations detailed in Section 3. We employ PoCC (PoCC, [n. d.]) to extract compile-time information such as schedules, loop trip counts, dependencies, and operation counts per statement. ISCC (Verdoolaege, 2011) then generates all legal permutations for each loop body.
Table 4 delineates the sets, variables, and constants utilized in our NLP formulation.
Constants Description
$II_{l}$ II of the loop $l$
$IL_{par}$, $IL_{red}$ Iteration Latency of the operations without ($par$) and with ($red$) dependencies of the statement $s$
$TC^{l}_{ori}$ Original Trip Count of the loop $l$
$f_{a,l}$ Footprint of the array $a$ if transferred to on-chip after the loop $l$
$N_{FT}$ Number of fused task
$DSP_{s_{op}}$ Number of DSP used by the statement $s$ for the operation $op$
$DSP$ Number of DSP available for the FPGA used
$max_{part}$ Maximum array partitioning
$SLR$ Number of SLR available for the FPGA used
Variables Description
$TC^{l}_{intra}$, $TC^{l}_{inter}$ TC of the loop $l$ for the intra and inter tile
$TC^{l}$ Trip Count of the loop $l$ after padding
$S^{last}_{a}$ Size of the last dimension of the array $a$ transferred on-chip
$BW_{a}$ Bit width of the array $a$
$t_{a,l},d_{a,l}$ Boolean to know if the array $a$ is transferred and defined (respectively) on-chip after the loop $l$ in the inter-tile
$p_{i}^{l}$ Position of the loop $l$ under the $i$-th permutation
$slr_{t}$ ID of the SLR use by the task $t$
Sets Description
$\mathcal{L},\mathcal{A},\mathcal{S}$ The set of loops, arrays and statements
$\mathcal{L}_{s}$ The set of loops which iterate the statement $s$
$\mathcal{L}_{s}^{red}$ The set of reduction loops which iterate the statement $s$
$\mathcal{L}_{a}^{last}$ The set of loop which iterate the last dimension of the array $a$
$\mathcal{L}_{inter},$ $\mathcal{L}_{intra}$ The set of loops which belong to the inter-tile and intra-tile respectively
$\mathcal{B}$ Set of possible burst size for the data type
$\mathcal{C}_{a_{d}}$ The set of loops which iterates the array $a$ at the dimension $d$
$AP_{a,d}$ Array Partition for the array $a$ in dimension $d$
$\mathcal{P}_{s}$ All permutation of the loops which iterate the statement $s$
$\mathcal{F}_{i}$ The set of statements which belong to the fused task $i$
$\mathcal{T}$ The set of tasks in the dataflow
Table 4. Mathematical Notation for Constants, Variables, and Sets in the NLP-Based Optimization Model. This table defines the formal notation used in Prometheus’ Non-Linear Programming (NLP) model for design space exploration, including loop trip counts, bitwidths, SLR mappings, resource limits, and legal transformations.
4.1. Constraints
We now describe the constraints by using the code of Listings 3, 4 and 5.
1for (int j1 = 0; j1 < 32; j1++) {
2#pragma HLS unroll
3 for (int i1 = 0; i1 < 19; i1++) {
4#pragma HLS unroll
5 for (int k1 = 0; k1 < 3; k1++) {
6#pragma HLS unroll
7 j=j0*32+j1; i=i0*19+i1; k=k0*3+k1;
8 F[i1][j1] += C[i][k] * D[k][j1];}
Listing 5: Implementation of Task 3 in the 3mm Kernel (from Listing 4). This listing illustrates the structure of Task 3, where the intra-tile computation is fully unrolled to expose fine-grained parallelism. The unrolling enables concurrent execution of loop iterations.
4.1.1. Data-tiling and Unroll Factor
The intra-tile transformation, as explained in Section 3, can divide either the original loop trip count or the original trip count with padding, thereby increasing the range of possibilities. Equation 1 ensures that the trip count of the intra-tile is a divisor of one of these two possibilities. The user has the option to constrain the padding using Equation 2, which simplifies the solution space for the NLP solver. For instance, in Listing 5, for the array $F$, the loop $j$ (line 7 in Listing 3) has been split into $j0$ (line 14 in Listing 4) and $j1$ (intra-tile). The trip count of the intra-tile loop $j1$, denoted as $TC^{j1}_{\text{intra}}=32$, does not evenly divide the original trip count $TC^{j}_{\text{ori}}=210$, but it does divide the trip count of the padded loop $TC^{j}=224$.
(1) $\forall l\in\mathcal{L},TC^{l}_{intra}\%TC^{l}_{ori}==0||TC^{l}_{intra}\%TC^{l}==0$
(2) $(opt)\forall l\in\mathcal{L},\exists n\in\mathbb{N}\leq N\in\mathbb{N},\text{ s.t. }TC^{l}=TC^{l}_{ori}+n$
4.1.2. Bit Width
$\mathcal{B}$ denotes the number of elements that can be transferred simultaneously, determined by the bit width and the data type. Hence, if we have a bit width under 512 bits for float the set is $\{1,2,4,8,16\}$. Equation 3 computes the bit width for each array based on the last dimension of the data-tile transferred on-chip. For example, the array D in Fused task 1 is transferred in line 15 with a size of $222\times 32$, so $S^{last}_{D}=32$, which is divisible by 16. Therefore, it has a bit width of 16.
(3) $\forall a\in\mathcal{A},\forall l\in\mathcal{L}_{a}^{last},\max_{b\in\mathcal{B}}BW_{a}=b\text{ s.t. }S^{last}_{a}\%b=0$
4.1.3. Permutation
Equation 4 requires the NLP to choose identical permutations for loops that are shared by statements fused within the same task. For example, in Fused task 1, the loops iterating statement S2 can be permuted as ($i0$, $j0$) or ($j0$, $i0$), and similarly, loops iterating statement S3 can be permuted as ($i0$, $j0$) or ($j0$, $i0$). However, since the loops $i0$ and $j0$ iterate both S2 and S3, they must use the same permutation; either ($i0$, $j0$) for both or ($j0$, $i0$) for both.
$\forall i\in\llbracket 0,N_{FT}\llbracket,\forall(ft_{0},ft_{1})\in\mathcal{F}_{i}^{2},$
(4) $\forall(i_{0},i_{1})\in\mathcal{P}_{ft_{0}}\times\mathcal{P}_{ft_{1}},\forall l\in\mathcal{L},p_{i0}^{l}=p_{i1}^{l}$
4.1.4. Transfer and Reuse
Equation 5 permits the selection of a single level where each array can be defined (and reused) and transferred. Equation 6 constrains that the definition of the array must occur lexicographically before or at the same time as the transfer.
The array $E$, defined on line 24 in Listing 4, is defined before any loops, so $d_{E,0}=1$ (0 indicating it is defined before any loops). However, it is transferred under the loop $i0$, so $t_{E,i0}=1$. Equation 6 simply means that the definition of array E should occur before or at the same level as the transfer. We cannot transfer E under loop $i0$ if E is defined under $k0$. Similarly, in Listing 4, the array $A$ is defined and transferred in line 4, with $d_{A,i0}=1$ and $t_{A,i0}=1$.
(5) $\forall a\in\mathcal{A},\sum l\in\mathcal{L}_{inter},t_{a,l}=1,\sum l\in\mathcal{L}_{inter},d_{a,l}=1$
(6) $\forall a\in\mathcal{A},\sum(l_{0},l_{1})\in\mathcal{L}_{inter}^{2}|d_{a,l_{0}},t_{a,l_{1}}=1,\text{ then }l_{0}\preccurlyeq l_{1}$
4.1.5. On-chip Memory
Equation 7 constrains the footprint of the array to be within the available resources, based on where the array is defined, the number of double buffers used and the footprint of the array transferred at this level.
(7) $\sum_{a\in\mathcal{A}}\sum_{l\in\mathcal{L}}d_{a,l}\times f_{a,l}\times N_{a}\leq Mem,$
with $N_{a}$ being the number of double buffer for the array $a$.
4.1.6. Array Partitioning
Equation 8 limits the maximum partitioning of each array. This partitioning is crucial as it impacts the maximum unroll factor, necessitating the distribution of data across different BRAM banks under fully unrolled loops, thereby influencing the utilization of BRAMs. Equation 9 computes the array partitioning needed for each array based on the trip count of the fully unrolled intra-tile loops.
Array D in Listing 5 is traversed by two unrolled loops: $k1$, which iterates 3 times ($AP_{D,0}=3$), and $j1$, which iterates 32 times ($AP_{D,1}=32$). Therefore, the total number of partitions needed is $3\times 32=96$. Consequently, these 96 values of F are stored in different banks, allowing all of them to be accessed in parallel. However, this value must be less than or equal to $max_{part}$.
(8) $\forall a\in\mathcal{A},\prod_{d\in\mathbb{N}}AP_{a,d}\leq max_{part}$
(9) $\forall a\in\mathcal{A},\forall d\in\mathbb{N},\forall l\in C_{a_{d}},AP_{a,d}=TC^{l}_{intra}==0$
4.1.7. DSP Utilization
Equation 10 constrains the number of DSPs used based on the available DSP resources. In contrast to (Pouget et al., 2025a, 2024, b), we utilize pessimistic DSP utilization. This approach is necessary because concurrent execution and resource reuse between tasks are not feasible when two tasks can run simultaneously. Given $DSP_{+}=2$, $DSP_{*}=3$, and $II_{S3}=3$, the DSP usage for Task 3 is calculated as $(2+3)\times 1824$, accounting for the unroll factor. However, since the loop is pipelined with $II=3$, the HLS compiler optimizes resource usage, effectively reducing the DSP count by approximately dividing by $II$.
(10) $\sum_{op\in\{+,-,*,/\}}\sum_{s\in\mathcal{S}}(DSP_{s_{op}}/II_{s})\times\prod_{l\in\mathcal{L_{s}}}TC^{l}_{intra}\leq DSP$
4.1.8. SLR Selection
For each task, the NLP determines the SLR on which the task will be implemented (Equation 11). Additionally, Equations 7 and 10 are applied to each SLR to manage resource allocation per SLR effectively.
(11) $\forall{t\in\mathcal{T}},slr_{t}\in\llbracket 0,SLR\llbracket$
4.2. Objective Function
The objective function models the dependency graph of the fused task. Independent tasks execute simultaneously, with latency set by the longest task. Dependent tasks overlap partially, with the second task starting once it receives enough data from the first. The latency is the maximum of both tasks’ durations plus the “shift”, the interval before the second task starts. For our 3mm example, the graph is shown in Figure 2, and the corresponding latency is:
$\left.\begin{cases}&Lat(T_{0},T_{1})=\max(Lat(T_{0}),Lat(T_{1})+shift_{T_{0},T_{1}})\\ &Lat(T_{2},T_{3})=\max(Lat(T_{2}),Lat(T_{3})+shift_{T_{2},T_{3}})\\ &Lat(T_{0},..,T_{4})=\max(Lat(T_{0},T_{1}),Lat(T_{2},T_{3}),Lat(T_{4}))\\ &Lat=\max(Lat(T_{0},..,T_{4}),Lat(T_{5})\\ &\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;+\max(shift_{T_{1},T_{5}},shift_{T_{3},T_{5}},shift_{T_{4},T_{5}}))\\ \end{cases}\right.$
PrometheusT0T1T2T3T4T5
Figure 2. Maximally distributed dataflow graph of the 3mm kernel, where each node represents a task and each edge indicates a data dependency between tasks.
The latency of the intra-task for a statement $s$ is determine similarly to (Pouget et al., 2025a, 2024, b) with Equation 12. And if the intra-task is iterated by the reduction loops in the inter-tile the latency is represented by Equation 13.
(12) $Lat_{intra}=IL_{par}+IL_{seq}\times\log_{2}(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{intra})$
(13) $Lat=Lat_{intra}+II\times(\prod_{l\in\mathcal{L}^{red}_{s}}TC^{l}_{inter}-1)$
For each level $n\geq 0$ (with the outermost level being 0) of the inter-tile, we consider the shifting of load, compute, and store operations to enable overlapping of computation and communication. The overlap factor $\alpha$ is set to 1 for read-only operations and 2 for read and store operations:
(14) $Lat_{n}=\max(Lat_{n+1},\frac{f_{a,l}}{BS_{a}})+Lat_{n+1}+\frac{f_{a,l}}{BS_{a}}\times\alpha$
And then, the final objective function is computed using the dataflow graph, where the latency of the task is established by the latency of the outermost level, denoted as $Lat_{0}$.
5. Code Generation
Prometheus takes as input affine C/C++ code and automatically produces an HLS-C++ file, OpenCL host files, and all necessary files for code verification, RTL simulation, and bitstream generation. The NLP described in Section 4 gives the parameters of the space such as loop order, tiling factor, etc. However, in order to be efficient, the code generation needs some specific rules.
5.1. Communication with Off-Chip Memory
To enable overlapping communication and computation, a load function transfers data on-chip using FIFOs, ensuring effective overlap management by the compiler. Furthermore, to guarantee that the load operation transfers data with the correct bit width, we automatically restructure the data in off-chip memory to enable sequential loading. Once data accumulates in the FIFO, a read function moves it to the shared data-tile buffer, whose size is determined by the NLP. For outputs, a write function transfers data from the buffer to a FIFO, and a store function writes it back to off-chip memory.
5.2. Communication between the Fused Task
The communication within the same fused task is not required as they use shared buffers. Similarly, for communication with off-chip memory, we choose to use FIFOs to facilitate communication between fused tasks, thereby simplifying the overlap.
5.3. Intra-Task
Each intra-task, corresponding to an intra-tile selected by the NLP, is implemented as a fully unrolled, independent function without communication with off-chip memory. Data for these tasks resides entirely on-chip.
If the task involves reduction loops, inter-tile reduction loops are integrated into the function and pipelined. While the pipeline initiation interval (II) is greater than 1 due to reduction dependencies, other operations in the statement are pipelined efficiently.
Padding is handled at the intra-tile level. Non-reduction loops remain unchanged, allowing computation of padding values without excessive resource usage. For reduction loops, full tiles are computed first, and the intra-tile loop is adjusted to handle padding for partial tiles accurately.
5.4. Inter-Tile Loop
For each fused task and each inter-tile loop, we generate independent functions to facilitate efficient double or triple buffering. Using information from the NLP, we determine which arrays are defined and which are transferred at each level of granularity. Once we identify the data being transferred, we implement double buffering if only reads or writes are involved, or triple buffering if both reads and writes are required. This strategy allows us to overlap the operations of reading, storing, and computing at the innermost level, optimizing both communication and computation overlap.
5.5. Concurrent Execution
Using the dataflow pragma, independent tasks execute concurrently without requiring manual rewrites. Computation in a receiving task begins as soon as its shared buffer contains sufficient data.
5.6. SLR Management
The NLP determines the SLR ID for each task, specifying where it will be executed. Prometheus generates a separate C++ file for each SLR, and data transfers between SLRs are managed via ap_axiu streams, ensuring efficient communication and minimizing transfer overhead.
5.7. Design Regeneration
In cases where bitstream generation fails for a given design, we support automatic regeneration of the design. Several strategies can be applied, such as tightening resource constraints or reducing the maximum unrolling factor. By reducing the available resources, the design becomes smaller, which can help alleviate congestion. Thanks to the flexibility of our NLP-based approach, we can retain parts of the previous solution—such as the SLR assignment—and selectively restrict resources only for the specific task or group of tasks responsible for the congestion. This is the method we employ to regenerate designs when bitstream generation fails.
6. Evaluation
6.1. Setup
We evaluated our method using kernels from Polybench/C 4.2.1 (Pouchet and Yuki, 2024) with medium-sized datasets and single-precision floating-point computations. The selected kernels represent both memory-intensive and computation-intensive scenarios. Medium problem sizes were chosen to balance demonstration of efficacy and the feasibility of time-consuming RTL evaluations. Due to Allo’s lack of automatic code generation, we limited our experiments to the subset of PolyBench kernels provided in its artifact evaluation package (all, 2024), as results for other kernels were not available. NLP problems were solved using the AMPL description language and the Gurobi solver (version 11.0.0) with the qp:nonconvex=2 option for non-convex quadratic objectives and constraints. Evaluations included RTL simulation and on-board execution using the Alveo U55C FPGA, with a targeted frequency of 220 MHz. RTL simulation provided accurate latency estimates, contrasting with the overly optimistic Vitis HLS reports that assume perfect task overlapping with dataflow pragma. The generated code from the frameworks are compiled using AMD/Xilinx Vitis HLS 2023.2 using the Vitis flow (AMD/Xilinx, 2024). This flow assumes that data initially resides off-chip and has a default latency of 64 cycles to bring onto on-chip memory. All frameworks utilize ”unsafe math” optimizations, enabling commutative/associative reduction operations at the expense of precision.
6.2. Experimental Evaluation
The objective of this evaluation is to demonstrate the capability of our framework to generate code with high QoR, whether for memory-bound or computation-bound kernels. To achieve this, we compare our work with Allo, ScaleHLS, Sisyphus, AutoDSE and Stream-HLS.
For ScaleHLS (Ye et al., 2022), Allo (Chen et al., 2024), and Stream-HLS (Basalama and Cong, 2025), their kernels assume that data is already present in on-chip memory. To ensure a fair comparison, we modified their code to incorporate off-chip to on-chip data transfers. However, Allo’s 2mm and 3mm kernels already include this transfer mechanism, requiring no modifications. We left Sisyphus and AutoDSE unchanged, as they already optimize bit width according to the problem size. We use AutoDSE with the bottleneck method, setting a DSE timeout of 1,000 minutes and an HLS synthesis timeout of 180 minutes per task. For Allo-generated designs, we utilize kernels from their artifact repository since Allo does not employ a DSE (all, 2024). Sisyphus generated designs are generated using parameters consistent with the paper (Pouget et al., 2025b). We observe some differences in Sisyphus results, as our evaluation is conducted on a different board. The design variations between boards may lead to differences in performance.
For RTL simulation, we assume that the frameworks can utilize all the resources on the U55C FPGA with a constraint of partitioning any array of 1,024 due to AMD/Xilinx limitations. For on-board FPGA evaluations, we consider two scenarios. The first scenario utilizes 60% of one Super Logic Region (SLR), equivalent to 20% of total board resources, for all frameworks. The second scenario is unique to our framework, leveraging all three SLRs, with 60% utilization per SLR. Most frameworks are not place-and-route aware, leading to congestion issues that prevent bitstream generation. Limiting frameworks to one SLR increases the likelihood of meeting timing requirements. AutoDSE, for instance, lacks dataflow support and applies pragmas within a single function, making multi-SLR bitstream generation impossible without manual intervention to split the function across SLRs. If congestion persists within an SLR, we adjust the NLP constraints to address the issue and regenerate the HLS-C++ file in just a few minutes.
It is worth noting that there is a difference in results between RTL simulation and on-board evaluation. This difference is expected, as RTL simulation involves fewer constraints, allowing for more aggressive unrolling and full utilization of available resources. In contrast, on-board evaluation is more complex, and excessive unrolling can easily lead to congestion. Therefore, we compare RTL simulation results with those of other frameworks to demonstrate our performance improvements. More importantly, we show that our approach generates code that can actually be implemented on board.
6.3. Comparison
Table 5 presents the RTL simulation results for Prometheus, Sisyphus (Pouget et al., 2025b), AutoDSE (Sohrabizadeh et al., 2021), ScaleHLS (Ye et al., 2022), and Allo (Chen et al., 2024). The last lines show the average and geometric mean performance improvement (PI) of Prometheus across evaluated kernels.
Kernel Ours Sisyphus ScaleHLS Allo AutoDSE Stream-HLS
2mm 308.38 195.09 37.13 46.58 0.41* 150.27
3mm 368.36 178.97 43.04 60.40* 1.74* 174.75
Atax 3.56 2.32 1.58 1.96 1.97* 1.71
Bicg 15.41 2.32 1.70 14.17 0.99* 1.73
Gemm 419.14 227.09 40.53 37.50 110.81* 203.48
Gesummv 10.21 2.28 1.78 8.85 1.98* 1.72
Mvt 14.65 5.54 7.39 8.77 7.80* 13.31
Symm 212.20 200.30 0.06 16.72 14.68* N/A
Syr2k 267.31 149.89 0.08 41.15 12.35* N/A
Syrk 158.25 105.59 0.27 20.57 23.16* N/A
Trmm 193.47 166.72 0.07 5.10 0.02* N/A
PI (Avg) 2.39x 927.20x 8.58x 973.14 3.46x
PI (gmean) 2.03x 48.03x 4.92x 25.82 2.71x
Table 5. Throughput Comparison (in GF/s) of PolyBench Kernels Across Frameworks Using RTL Simulation
Results marked with * are from the Vitis HLS report, showing minimum latency as an optimistic estimate. The RTL simulation for 3mm with Allo was incomplete after two days. The N/A values in the table for Scale-HLS correspond to kernels that have at least one loop with a non-constant trip count. Stream-HLS does not handle these cases, so we cannot make a comparison.
Prometheus consistently achieves superior QoR across all evaluated kernels compared to other frameworks. While Sisyphus (Pouget et al., 2025b), designed primarily for computation-bound kernels, demonstrates competitive results for these kernels, it shows a weakness for 2mm and 3mm. This disparity is due to Sisyphus lacking concurrent execution capabilities for independent matrix multiplications. The performance advantage of Prometheus stems from both concurrent task execution and efficient overlapping of computation and communication.
Although for bicg, Allo and Prometheus do not use the same code transformation, the results are similar. Allo retains the original code structure by permuting the reduction loop outermost. The non-reduction loop is fully unrolled, and the reduction loop is pipelined. Prometheus partially unrolls both loops and pipelines the reduction loop.
On Board Evaluation  Table 6 presents the results for the on-board evaluation. The column $T(ms)$ indicates the kernel execution time in milliseconds, GF/s represents the throughput in Giga Floating Operations per second, and the resource usage is detailed in the thousands for both LUTs (L) and FFs. URAM utilization is excluded as no kernels use it. Column F (MHz) shows the frequency achieved by each design, with a target frequency of 220 MHz for all designs.
For Bicg and Atax, targeting 60% of resources on one SLR caused congestion, requiring regeneration with a 55% constraint. The 3mm bitstream from AutoDSE succeeded only with a 15% constraint.
Kernel T (ms) GF/s DSP BRAM L(K) FF(K) F (MHz)
1 SLR Sisyphus 2mm 1.20 30.57 556 510 213 276 220
3mm 1.52 29.89 984 611 230 300 220
Atax 0.62 1.03 173 450 240 250 220
Bicg 0.63 1.02 173 451 238 265 217
1 SLR AutoDSE 2mm 92.25 0.40 963 353.5 287 292 205
3mm 110.34 0.41 1117 470 278 306 220
Atax 2.88 0.22 452 630.5 170 212 220
Bicg 1.13 0.56 196 867.5 168 217 214
1 SLR Ours 2mm 0.56 65.13 1941 635.5 371 454 216
3mm 0.87 51.95 1551 635.5 342 423 220
Atax 0.24 2.62 1081 533.5 234 287 184
Bicg 0.15 4.04 732 311.5 250 302 220
3 SLR Ours 2mm 0.29 125.54 2752 546 428 549 220
3mm 0.34 134.07 4379 600 684 840 207
Atax 0.20 3.10 1823 634.5 405 539 137
Bicg 0.14 4.34 1226 241 291 380 177
Table 6. On-Board Evaluation of Performance and Resource Utilization Across Frameworks and SLR Configurations
Prometheus achieves a remarkable 77.16x performance improvement over AutoDSE. This gain is accompanied by an average increase in resource utilization: 2.38x more DSPs, 1.08x more BRAM, 1.36x more LUTs, and 1.42x more FFs, reflecting the trade-offs made to achieve such high efficiency. When compared to Sisyphus, Prometheus demonstrates a notable 2.59x performance boost, leveraging an average of 3.88x more DSPs, 1.04x more BRAM, 1.31x more LUTs, and 1.33x more FFs, illustrating its ability to deliver substantial improvements while effectively utilizing additional resources.
Table 7 shows the parameters found by the NLP. We use the same name of the iterator as Polybench 4.2.1 (Pouchet and Yuki, 2024) code. $Si$ represent the statement in position $i$ in the code. The second column gives the statement fused inside the same tasks, the third column sets the fused task order and loop order found by the NLP for the fused task and the last column supplies the data-tile size found by the NLP, if the array is present in a different fused task the fused-task (defined in the second column) is precise.
Permutations are evident in the implementations of 3mm, Atax, and Bicg. For 2mm and 3mm, the NLP opted to fully transfer array B instead of overlapping computation and load, as the load operation had higher latency than computation. In 2mm, the NLP retains the original loop order. Array tmp is transferred from the first task to the second, with both tasks iterating over the first dimension using loop $i$. This enables a $4\times 32$ data tile to be sent to the second task, which starts computation once a $4\times 192$ tile of tmp(FT1) is filled.
Fused Statement Loop   Order Data Tile Size
2mm FT0: S0, S1, FT1: S2, S3 FT0: i,j,k, FT1: i,j,k tmp(FT0): $4\times 32$, B: $212\times 192$, A: $4\times 212$, D: $4\times 32$, C:$192\times 32$, tmp(FT1): $4\times 192$
3mm FT0: S0, S1, FT1: S2, S3, FT2: S4, S5 FT0: i,j,k, FT1: j,i,k FT2: j,i,k E(FT0): $9\times 16$, A: $9\times 200$, B: $200\times 192$, F(FT1): $10\times 10$, C: $10\times 224$, D: $224\times 10$, G: $9\times 10$, E(FT2): $9\times 192$, F(FT2): $192\times 10$,
Atax FT0: S1, S2, FT1: S0, S3 FT0: i,j, FT1: j,i tmp(FT0): $56$, y: $16$, A(FT0): $392\times 416$, A(FT1): $392\times 16$, tmp(FT1): $392$
Bicg FT0: S1, S2, FT1: S0, S3 FT0: j,i FT1: i,j s: $16$, A (FT0): $416\times 16$, r: $416$, q: $10$, A (FT1): $10\times 400$, p: $400$,
Table 7. Fusion, loop order and data-tile size found by the NLP for the kernel evaluated on Table 6 for 1 SLR
As other frameworks cannot generate bitstreams for 3 SLRs without human intervention, we compare results for 1 and 3 SLRs using our framework. For 2mm and 3mm, performance improves due to increased resource utilization. However, for atax and bicg, where the bottleneck is memory transfer between off-chip and on-chip rather than parallelism, the improvement is negligible.
6.4. Scalability
Our NLP solver includes the option to set a timeout, returning the best design found so far without guaranteeing optimality. This feature enables efficient exploration of the solution space while ensuring adherence to time constraints when necessary.
When analyzing the time required to solve the Non-Linear Programming (NLP) problem in Sisyphus, we observe that our framework, Prometheus, achieves similar solution times for most benchmarks. However, there is a notable exception: the 3mm benchmark, which times out after 4 hours in Sisyphus, while Prometheus successfully finds a solution.
This efficiency stems from Prometheus’ ability to explore a larger optimization space while still maintaining fast solution times. The key reason is that all optimization techniques are seamlessly integrated within the design space. This structured integration ensures that when a decision is made for one optimization parameter, it naturally constrains the possible choices for others, thereby reducing the search complexity. As a result, even with a broader search space, Prometheus efficiently converges to a solution in 21.37 seconds.
A detailed comparison of solution times can be found in Table 8, which highlights the impact of our approach in managing the design space effectively.
Benchmark Sisyphus Prometheus
2mm 22.37 15.98
3mm 14400.08 21.37
Atax 3.18 7.03
Bicg 3.17 5.09
Gemm 2.19 4.31
Gesummv 1.57 8.07
Mvt 1.71 1.26
Symm 7.13 11.46
Syr2k 5.29 6.07
Syrk 2.56 3.94
Trmm 4.34 5.96
Average 1313.96 8.23
Geo Mean 7.95 6.50
Table 8. Time Required (in seconds) for the NLP Solver to Find an Optimal Solution: Sisyphus vs. Prometheus Across Benchmarks (Timeout Set at 14,400s)
7. Related Work
Dataflow  Dataflow principles have been extensively studied in models such as Kahn Process Networks (Kahn, 1974), Dennis Dataflow (Dennis, 1974), synchronous dataflow languages (Lee and Messerschmitt, 1987; Benveniste et al., 1994), and for FPGA applications (Peverelli et al., 2018; Bhattacharyya et al., 2009; Abid et al., 2013; Grigoraş et al., 2013; Ben-Nun et al., 2019; Amiri et al., 2021; Xiang et al., 2022; Chen et al., 2024). DaCe (Ben-Nun et al., 2019) introduces Stateful DataFlow multiGraphs to separate program definition from optimization, enabling transformations like tiling and double-buffering, though requiring user intervention. Stream-HLS (Basalama and Cong, 2025) automatically generates dataflow; however, it assumes that data are already on-chip, thereby overlooking communication with off-chip memory. Additionally, it offers very limited opportunities for parallelism. Flower (Amiri et al., 2021) automates FPGA dataflow development but limits parallelism. Frameworks like (Chen et al., 2024; Xiang et al., 2022), built on HeteroCL (Lai et al., 2019), optimize data placement and compute scheduling for heterogeneous systems, maximizing data reuse and bandwidth utilization. Systolic arrays (Wang et al., 2021; Cong and Wang, 2018; Jia et al., 2020; Lai et al., 2020; Basalama et al., 2023) offer efficient computation for specific patterns but lack generalization. Application-specific frameworks (Chi et al., 2018; Hegarty et al., 2014; Natale et al., 2016; de Fine Licht et al., 2021; Singh et al., 2020; Denzler et al., 2023; Korol et al., 2022; Bacis et al., 2017) demonstrate dataflow advantages but do not generalize across domains. RapidStream-TAPA (Guo et al., 2023a, b) enhances the performance of dataflow designs and automates SLR placement. However, it requires an optimized kernel with a dataflow structure as input.
Shared Buffer  Shared buffer utilization through HLS has been extensively explored using methods such as NLP-based pragma insertion (Pouget et al., 2024, 2025a, 2025b), bottleneck DSE (Sohrabizadeh et al., 2021), and GNN-based latency and resource estimation (Sohrabizadeh et al., 2023, 2022a, 2022b; Bai et al., 2022; Ustun et al., 2020; Wu et al., 2022a, b). However, these approaches lack effective integration of dataflow optimization techniques.
Code Transformation  Code transformation has been explored for CPUs (Bondhugula et al., 2008; Pouchet et al., 2011; Kruse et al., 2020; Baghdadi et al., 2019), GPUs (Verdoolaege et al., 2013), and FPGAs (Pouchet et al., 2013b; Zhao et al., 2022; Zhao and Cheng, 2021; Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018). Pluto (Bondhugula et al., 2008) minimizes communication and improves locality but can limit parallelism. FPGA-specific adaptations (Pouchet et al., 2013b) leverage FIFOs for overlapping communication and computation but are restricted in parallelism. Recent works (Zhao et al., 2022; Zhao and Cheng, 2021) selectively use Pluto for latency minimization while avoiding non-HLS-friendly code. While (Li et al., 2014; Liu et al., 2016, 2017b, 2015; Choi and Cong, 2018) focus on optimizing pipelining techniques, they do not address parallelism or the coordination of computation and communication overlap, which are crucial for our objectives. The (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021; Chen et al., 2024; Ye et al., 2024) compilers perform code transformations and pragma insertion, their modifications are primarily heuristic and based on loop properties. The paper (Pouget et al., 2025b) described in Section 2 generates design with high QoR, but the absence of dataflow utilization hinders concurrent task execution. Additionally, their approach avoids padding, limiting the unroll factor to divisors of the loop’s trip count and constraining tiling space.
Tiling and Padding  Tiling is essential for balancing computation and communication. While prior works (Liu et al., 2017a; Li et al., 2021) use cost models to minimize communication, our approach extends this to reduce overall latency. Techniques like NLP-based tiling (Li et al., 2021) focus on CPUs, while Wedler (Shi et al., 2023) optimizes DNNs on GPUs by fusing operators, enhancing data reuse, and employing padding to prevent bank conflicts. Padding is well-studied for reducing cache misses (Hong et al., 2016; Panda et al., 1999) and improving memory transfers (Wang et al., 2012), but its use for varying unroll factors on FPGAs remains underexplored.
8. Conclusion
In this work, we introduced Prometheus, a holistic optimization framework that unifies loop transformations, task concurrency, computation-communication overlap, and hardware-aware scheduling for FPGA accelerators. By formulating the optimization process as a Non-Linear Programming (NLP) problem, Prometheus enables a structured exploration of the vast design space while considering hardware constraints, memory bandwidth, and task parallelism.
Our framework addresses key limitations in existing methodologies, which often optimize only isolated aspects of FPGA acceleration. Prometheus surpasses these approaches by integrating SLR-aware scheduling, dynamic memory management, and hybrid execution models that effectively balance dataflow streaming and shared buffering. This enables more efficient utilization of FPGA resources, leading to significant improvements in latency, throughput, and scalability.
Through extensive performance evaluations on computation-bound kernels, we demonstrated that Prometheus outperforms state-of-the-art frameworks.
Furthermore, Prometheus’ automatic design space exploration significantly reduces the manual effort required in FPGA development. By automatically generating HLS-C++ code, OpenCL host code, and FPGA bitstreams, the framework streamlines the deployment process, making FPGA acceleration more accessible to a broader range of applications.
In summary, Prometheus provides an innovative and effective approach to FPGA optimization, delivering high-performance solutions while minimizing the complexity of hardware design.
Acknowledgments
This work was supported by the NSF award #CCF-2211557. It is also supported by CDSC industrial partners and the AMD/HACC Program.